### v1

# Bayesian sample size calculations for external validation studies of risk prediction models

Abstract
Background: Contemporary sample size calculations for external validation of risk prediction models require users to specify fixed (i.e., true) values of assumed model performance metrics alongside target precision levels (e.g., 95% CI widths). However, due to the finite samples of previous studies, our knowledge of true model performance in the target population is uncertain, and so choosing fixed values represents an incomplete picture. As well, for net benefit (NB) as a measure of clinical utility, the relevance of conventional precision-based inference is doubtful. Bayesian approach can address this, by accounting for uncertainty of true model performance and facilitating decision-theoretic metrics including the expected value of sample information (EVSI).
Methods: We propose a general Bayesian algorithm for constructing the joint distribution of predicted risks and response values based on summary statistics of model performance in previous studies. For statistical metrics of performance, we propose sample size determination rules that either target desired expected precision, or a desired assurance probability that the precision criteria will be satisfied. For NB, we propose rules based on optimality assurance (the probability that the planned study correctly identifies the most beneficial strategy) and the EVSI, the expected gain in net benefit from the planned validation study. We showcase these developments in a case study on the validation of a risk prediction model for deterioration of hospitalized COVID-19 patients.
Results: The contemporary approach, based on fixed values of assumed c-statistic, O/E ratio, and calibration slope from in the validation study, with target 95%CI width of, respectively, 0.10, 0.22, and 0.30, would recommend a sample size of 1,056, dictated by the desired precision around the calibration slope. To implement the Bayesian approach and to account for the uncertainty of the model’s predictive performance, we constructed predictive distributions for prevalence, c-statistic, mean calibration error, and calibration slope for a new validation study. Targeting the same expected CI width would result in a sample size of 1,025. However, demanding 90% assurance for meeting these precision criteria would result in a sample size of 1,173. In terms of NB, an optimality assurance of 90% was achieved at a sample size of 306. The EVSI curve demonstrated a diminishing margin beyond sample sizes >500 (<10% gain in expected NB with doubling the sample size).
Conclusion: Compared to the conventional sample size calculation methods, a Bayesian approach requires explicit quantification of uncertainty around model performance, but thereby enables various sample size rules based on expected precision, assurance probabilities, and value of information. In our case study, EVSI calculations indicated that the precision criterion around calibration slope could potentially be relaxed without much utility loss.
1 Background
Developing risk prediction models or validating existing ones in a new population represents significant investment in time, resources, and expertise. Like other empirical experiments, the design of such studies should be based on objective, transparent, and defendable principles. A particular aspect of study design is the sample size of such studies. The field has witnessed significant recent developments on this front1–8. An example is the multi-criteria approach by Riley et al for the sample size required for external validation studies, targeting, targeting pre-specified width of the 95% confidence interval around metrics of model performance2,3. For binary outcomes, such metrics are related to discrimination, calibration, and net benefit (NB). The sample size required for each component is computed separately, with the largest one advertised as the final requirement3.
In addition to target precision criteria, each component of this approach requires as input an assumed true value of the the metric of interest in the target population. In reality, we do not know the true value of model performance metrics with certainty (otherwise there would be no need to conduct the study in the first place). Further, due to sampling variability, the precision obtained in one particular dataset may differ substantially from the expected precision for that sample size. Thus, there is more uncertainty to be accounted for than the Riley criteria allow.
A Bayesian approach towards sample size determination allows uncertainty to be accounted for, which is advantageous on multiple fronts. First, it enables the full use of existing information on model performance at the time the study is designed, rather than forcing the investigator to express their knowledge as the fixed truth. Second, it enables different classes of sample size rules, namely those that target the expected values of precision targets, as well as ‘assurance’-type rules for the probability of meeting (or exceeding) precision targets. This can be insightful as focusing on expected values alone does not guarantee that the desired precision will be achieved in one particular dataset, and the investigator may want a stronger assurance against this. For example, an investigator that desires an interval width of 0.1 for c-statistic might not perceive much benefit from the future interval being narrower, but might have a strong preference against not meeting this target. As such, targeting a sample size that will result in 90% probability that the CI width will be $\leq$ 0.1 might be more appealing than a sample size that targets an expected CI width of 0.1. Finally, when assessing clinical utility, the relevance of precision-based criteria is challenged9,10. A Bayesian approach enables the use of novel, decision-theoretic approaches based on value of information (VoI) analysis, which focus on the expected gain in clinical utility from increasing sample size11.
Bayesian approaches for power and sample size calculations are mostly investigated in the context of experimental studies that are aimed at interrogating a null hypothesis12–15. However, risk model validation studies are not generally hypothesis-driven. Rather, the focus is on the precise estimation of a variety of metrics of model performance. As such, Bayesian sample size considerations in this context is worth exploring. Hence, in this article, we propose a Bayesian version of the sample size formula by Riley et al, focusing on the same metrics of model performance, but enabling the investigator to 1) incorporate their uncertainty around assumed model performance in calculations, 2) use sample size rules that target assurance probabilities for meeting chosen criteria, and 3) use rules based on VoI analysis for clinical utility. The proposed framework can be used in two general ways: to quantify the anticipated precision or VoI outcomes from a planned study if the sample size is fixed (as is the case with validation studies that are based on already collected data), or to determine the minimum sample size that achieves pre-specified precision or VoI criteria. This framework can also be used in a hybrid form: determine the sample size based on certain criteria (e.g., targeting CI width for c-statistic and calibration slope), and investigate the consequence of the chosen sample size using other criteria (e.g, assurance probability around NB for the final sample size). Calculations can be done for the entire sample, as well as within subgroups, for example, imposing fairness criteria around the precision of estimates among minority sub-groups.
The rest of this manuscript is structured as follows. First, we briefly review common metrics of model performance and multi-criteria sample size formula by Riley et al. We will then introduce a Bayesian extension of this framework based on various sample size rules. We propose a general approach for characterizing uncertainty around model performance based on commonly reported information from previous studies. We outline Monte Carlo sampling algorithms for drawing from the posterior distributions of relevant quantities. A case study based on a model for predicting COVID-19 deterioration showcases the developments. We conclude by suggesting further areas for research.
2 Methods
We review the context, common metrics of model performance, Riley’s sample size formulas for external validation studies, and our proposal for Bayesian sample size determination.
2.1 Context
We focus on external validation of a model for predicting the risk of a binary outcome. Broadly speaking, an external validation study is an endeavor in learning about the joint distribution of predicted risks ($\pi$) from a pre-specified model and observed outcomes ($Y$: 0 no-event, 1 event) in a target population. The validation sample $D_{N}$ of size $N$ can be seen as $N$ pairs of predicted risks and observed results: $D_{N}=\{(\pi_{i},Y_{i})\}_{i=1}^{N}$.
A classical validation study does not involve learning about the relationship between predictors and the outcome. Rather, the focus is on quantifying the performance of a pre-specified model, and ultimately identifying whether it is worth implementing in the target population. For ease of expositions, we assume one model is being considered. This framework can be extended to multi-model comparisons with relative ease, but we leave this to subsequent exploration. As is the standard in contemporary practice, predicted risks for the model are taken as fixed values, representing the expected outcome probability among all individuals with the same predictor pattern. The proposed approach is applicable for both classical regression models and black-box machine learning algorithms, as a validation study is not concerned on how predictors are used to compute the predicted risk.
2.2 Common metrics of model performance
Most model development and validation studies report the following metrics along with associated uncertainty:
1) Outcome prevalence: $\phi:=\mathbb{E}(Y)$.
2) Calibration function $h()$: This is the function that returns the expected actual outcome risk among individuals with a given predicted risk: $h(\pi)=P(Y=1|\pi)$. This function converts the predicted risk for the ith individual, $\pi_{i}$, to the true risk for that person, which we denote by $p_{i}$. $p_{i}:=h(\pi_{i})$. When using non-functional form $h$ we refer to the parameters that define $h()$. Commonly, including in Riley’s sample size formula, $h()$ is modeled linearly on the logit scale: $\mbox{logit}(h(\pi))=\alpha+\beta\mbox{logit}(\pi)$, and thus $h$ consists of an intercept ($\alpha$) and slope ($\beta$). Sometimes studies report $\beta$ and ‘mean calibration’, i.e., $\mathbb{E}(Y-\pi)$, or $\beta$ and observed-to-expected outcome ratio ($O/E:=\mathbb{E}(Y)/\mathbb{E}(\pi)$), but these are equivalent as for a fixed value of $\beta$, knowing any of calibration intercept, mean calibration, or O/E ratio is enough to specify the calibration line. In other instances, $h()$ is modeled more flexibly using non-parametric methods (e.g., based on LOESS smoothing)16. Riley et al, while determining sample size based on a logit-linear calibration function, suggests visually inspecting the variability in anticipated smoothed curves once the sample size is determined2.
3) c-statistic $(c)$: This is a measure of the discriminatory performance of the model, and is the probability of concordance between the ranking of predicted risks and outcomes among a randomly chosen pair. Formally, $c:=P(\pi_{2}>\pi_{1}|Y_{2}=1,Y_{1}=0)$ with $(\pi_{1},Y_{1})$ and $(\pi_{2},Y_{2})$ being two randomly selected pairs of predicted risks and responses.
4) Net benefit (NB): Details of NB calculations are provided elsewhere17. In a nutshell, at a chosen risk threshold $z$, there are at least three treatment strategies concerning the use of the model: treat no one (with a default NB of 0), use the model to treat individuals whose predicted risk is $\pi\geq z$, or treat all. We index these three strategies by 0, 1, and 2, respectively. The $NB()$ function can be written as (for brevity of notations, in what follows we drop the notation that would indicate NB-related calculations are based on the chosen risk threshold):
$NB(k)=\left\{\begin{array}[]{lll}0&k=0&\mbox{(treat no one)}\\ \phi se-(1-\phi)(1-sp)\frac{z}{1-z}&k=1&(\mbox{use model to decide})\\ \phi-(1-\phi)\frac{z}{1-z}&k=2&(\mbox{treat all})\\ \end{array}\right.,$
where $se:=P(\pi\geq z|Y=1)$ and $sp:=P(\pi<z|Y=0)$ are the sensitivity and specificity of the model at the chosen threshold, respectively. An important difference between NB and other metrics of model performance is that NB is a decision-theoretic measure. If using the model has higher expected NB over other strategies (irrespective of uncertainties), then it is expected to confer clinical utility and so the decision would be to use the model. On the contrary, metrics of model performance such as c-statistic or calibration function do not consider the decision-making context.
2.3 Riley’s sample size formulas
Citing the commonality of the above-mentioned metrics in risk modeling studies, Riley et al structured their multi-criteria equations around desired precision levels (width of 95%CI) around these metrics3. In particular, the following equations were suggested for approximating the standard deviation of the sampling distribution of the estimator (i.e., standard error [SE]), from which Wald-type confidence intervals could be constructed:
$\mbox{SE}(c)=\sqrt{\frac{c(1-c)\left[1+(N/2-1)(\frac{1-c}{2-c})+\frac{(N/2-1)c}{1+c}\right]}{N^{2}\phi(1-\phi)}},$
$\mbox{SE}(\log(O/E))=\sqrt{\frac{1-\phi}{N\phi}},$
and
$\mbox{SE}(\beta)=\sqrt{\frac{I_{\alpha}}{N(I_{\alpha}I_{\beta}-I^{2}_{\alpha,\beta})}},$
where $I_{\alpha}={\mathbb{E}}\left(p(1-p)\right)$, $I_{\beta}={\mathbb{E}}\left(\mbox{logit}(\pi)^{2}p(1-p)\right)$, and $I_{\alpha,\beta}={\mathbb{E}}\left(\mbox{logit}(\pi)p(1-p)\right)$, and $p=h(\pi)$.
Assuming all measures are of interest, the final sample size is decided by the largest $N$ among each component. Riley et al proposed also targeting desired CI bands around standardized NB. However, a Bayesian approach facilitates using decision-theoretic approaches for NB, addressing the criticisms around the relevance of conventional inferential methods for NB9,10.
Note that the calculations require assumptions about the anticipated model performance (e.g., c-statistic, calibration slope) and the distribution of predicted risks in the target population. The authors suggested using single point-estimates for the true model performance values, and one chosen distribution for predicted risks. We now propose a Bayesian approach that enables modeling uncertainties around these quantities.
2.4 Going Bayesian
A Bayesian approach towards sample size calculation requires 1) identifying metrics of interest for model performance that will be estimated from this sample (e.g., c-statistic, calibration slope), 2) deciding on precision targets for such metrics (e.g., a 95% CI width of 0.1 for c-statistic), 3) deciding on sample size rules on such metrics (e.g., targeting the expected CI width, demanding a 90% assurance that the CI width will be met, or 90% assurance that the future validation study will correctly identify the most optimal decision at a risk threshold of interest; 4) specifying our current information about model performance; 5) generating random draws from the posterior distribution of desired targets that will be processed according to the specified rules.
In line with the contemporary practice in risk prediction development and validation, we adopt a mixed Bayesian-likelihood perspective: we use existing evidence (prior) to quantify our uncertainty about model performance, but we assume once the future sample is obtained, we will solely rely on it (the likelihood). We will discuss the implications of a fully Bayesian approach where the likelihood and prior are combined.
Let $P_{\theta}(\pi,Y)$ represent our knowledge about the joint distribution of predicted risks and outcomes in the target population. This joint distribution is indexed by parameters $\theta$, a random entity summarizing our current knowledge about this joint distribution (which encompasses our beliefs about model performance in the target population). We learn about $\theta$ by collecting ${D}_{N}$. The key aspect of a Bayesian approach is that we treat $\theta$, $D_{N}$, and therefore any quantity (e.g., CI widths) derived from $D_{N}$, as random entities. Implementing this approach involves the following steps: we specify $P(\theta)$, our current (prior) information about model performance in the target population. Repeatedly, we sample from $P(\theta)$ and simulate a random validation sample $P(D_{N}|\theta)$, from which performance metrics are quantified and recorded. Repeating this many times will generate draws from the posterior distribution of these performance metrics for a validation dataset of a particular sample size. These can be summarized in terms of their mean (expected value) and variance, and lead to various sample size rules and assurance probabilities.
Naturally, Bayesian Monte Carlo sampling results in draws from posterior distributions of targets (e.g., CI widths) given a fixed $N$. Sample size calculation thus becomes a stochastic inverse problem: find the smallest $N$ that satisfies a set of sample size rules. As the latter is detached from the core of a Bayesian workflow, we primarily focus on calculating precision and VoI metrics for a given $N$, and discuss our implementation of stochastic root-finding algorithms for sample size determination afterwards.
2.4.1 Sample size rules
Because the Bayesian approach takes estimates of model performance in the future validation study as random quantities, it invites establishing sample size rules that take into account such randomness. For statistical measures of model performance, we consider two sets of rules: those that target expected precision intervals, and those that target a probability (assurance) that the anticipated precision will be at least as good as targeted. For NB as a measure of clinical utility, we suggest sample size rules that are not precision-based and rather hone in on our ability in detecting the strategy with the highest NB.
2.4.1.1 Sample size rules for metrics of discrimination and calibration
Broadly speaking, we consider some scalar summary derived by applying summarizing function $g()$ to validation data ${D}_{N}$, to meet some criterion. In our context, $g({D}_{N})$ returns the 95%CI width for the corresponding metric.
Expected CI widths (eciw): This is related to the Average Length Criterion discussed by Joseph et al13. Here, we target the expected CI width (eciw) across the distribution of ${D}_{N}$:
$eciw(N)=\mathbb{E}_{\mathbf{\theta}}\mathbb{E}_{{D}_{N}|\mathbf{\theta}}(g({D}_{N}))$
If the sample size is fixed, $eciw(N)$ is reported for each component as the expected precision of the future study. For sample size calculation, we seek the minimum $N$ that results in $eciw(N)$, meeting precision target $\tau$: $\min\{N\in\mathbb{N}|eciw(N)\leq\tau\}$. We derive $N$s separately for each component, and choose the maximum $N$ as the final sample size. For example, we might identify the $N$ required to ensure that the expected CI widths of the calibration slope, and c-statistic all meet a desired expected CI width.
Assurance-type rules based on quantiles of CI width (qciw): This is related to the modified worst outcome criterion as discussed by Joseph et al13. Here, we target the probability of meeting (or exceeding) desired precision targets. For a given sample size, this approach returns the CI widths corresponding to a desired quantile $q$ (e.g., $q=0.9$ for 90% assurance):
$qciw(N,q)=F_{N}^{-1}(q),$
with
$F_{N}(q)=P(g(D_{N})\leq q)=\mathbb{E}_{\mathbf{\theta}}\mathbb{E}_{{D}_{N}|\mathbf{\theta}}(I(g({D}_{N})\leq q))$
being the CDF of the distribution of the anticipated CI width for sample size $N$.
Again, for a fixed-$N$ setup we report $qciw(N)$. For sample size determination, we find the minimum $N$ such the the $q^{th}$ quantile is not greater than the target CI width $\tau$: $\min\{N\in\mathbb{N}|qciw(N,q)\leq\tau\}$.
2.4.1.2 Sample size rules for NB
Optimality assurance for NB : This is the expected probability that we will correctly identify the strategy that has the highest population NB based on $D_{N}$. To proceed, let $NB_{D_{N}}(k)$ be the sample estimate of $NB(k)$, and $NB_{\theta}(k)$ its true value given $\theta$. We note that with the future data at hand, the investigator will declare a winning strategy as the one that has the highest expected NB solely based on the sample:
$W(D_{N})=\mbox{arg}\max_{k}(NB_{D_{N}}(k))$
Optimality assurance is the probability that the NB of the strategy that we will declare as the best in the sample is the maximum possible NB:
$Assurance(N)=P\left(NB_{\theta}(W(D_{N}))=\max_{k}(NB_{\theta}(k))\right)=\mathbb{E}_{\theta}\mathbb{E}_{D_{N}|\theta}I\left(NB_{\theta}(W(D_{N}))=\max_{k}(NB_{\theta}(k))\right).$
This assurance is in the probability scale that is non-decreasing and asymptotic to 1 as the sample size is increased.
Expected Value of Sample Information: The key VoI quantity is EVSI(N), the expected gain in NB from conducting a future validation study of size $N$.
$EVSI(N)=\mathbb{E}_{\theta}\mathbb{E}_{D_{N}|\theta}\left(NB_{\theta}(W(D_{N}))\right)-\max_{{}_{k}}{\mathbb{E}_{\mathbf{\theta}}NB_{\theta}(k)}.$
The first term on the right hand side is the expcted NB of the decision that we will declare as optimal based on $D_{N}$. The second term on the right-hand side is the expected NB of the best decision under current information. However, it might be the case that without the validation study, the model will not be implemented, regardless of its potential superiority under current information. In which case, this term can be replaced by the NB of the default strategy (e.g., treating no one or treating all).
EVSI(N) is a non-decreasing function. Its maximum value occurs when $N$ is infinity, that is, we have access to the entire population and can unequivocally determine the best strategy. The expected gain from such perfect information is called the expected value of perfect information (EVPI) and can be calculated as (18,19):
$EVPI=\mathbb{E}_{\mathbf{\theta}}\max_{k}(NB_{\theta}(k))-\max_{k}{\mathbb{E}_{\mathbf{\theta}}NB_{\theta}(k)}.$
EVSI and EVPI can be expressed in true positive or false positive units. As these units are context-specific, it is more natural to propose a unit-less metric as a target of sample size rule. We propose ‘relative’ EVSI (rEVSI) as the ratio of EVSI to EVPI. This value is intuitive and can be presented as percentage. An $rEVSI$ of 0.4 for a given sample size means that the external validation study at this sample size is expected to reduce the expected NB loss due to uncertainty by 40%.
Another sample size rule can be based on the expected net benefit of sampling (ENBS)20. ENBS requires scaling the EVSI by the expected size of the population affected by the decision, and subtracting the costs of collecting a sample in the same (true or false positive) unit as NB. Let $\mathcal{M}$ be the expected number of times the decision of interest is to be made, and $\mathcal{W}$ be the effort of every additional recruitment for the validation study in NB units. Then
$ENBS(N)=\mathcal{M}EVSI(N)-\mathcal{W}N,$
and the optimal sample size is the one that maximizes ENBS. This rule for sample size determination is fully decision-theoretic as it avoids specifying any arbitrary threshold values. However, it requires scaling total population size and establishing the trade-off between sampling efforts and clinical utility; both of these tasks are context-specific. As such, while we propose this rule here for completeness, we will not investigate it in the case study.
2.4.2 Specifying $P(\theta)$
The Bayesian approach requires specifying $P(\theta)$, the distribution of parameters that govern $P_{\theta}(\pi,Y)$. This parameterization can be done in different ways. For example, if a (pilot) sample form the target population is available, one can adopt a non-parametric approach based on the Bayesian bootstrap. In this scheme, $\theta$ is the vector of weights assigned to each observation in the pilot sample. There weights are random with a distribution of $P(\theta)\sim\mbox{Dirichlet}(1,...,1)$. The empirical distribution of this weighted sample can be considered as a random draw from the distribution of the population21. The validation sample $D_{N}$ then can be obtained from sampling with replacement. Details of this two-level resampling approach is provided elsewhere18,22.
Our focus here is on parametric modeling based on summary statistics from previous studies. Noting that $P(\pi,Y)=P(\pi)P(Y|\pi)$, with $Y|\pi\sim\mbox{Bernoulli}(h(\pi))$, an intuitive way of parameterizing $\theta$ is via specifying $P(\pi)$, the distribution of predicted risks, and $h$, the parameters defining the calibration function $h()$. However, specifying the distribution of predicted risks in the target population is not straightforward, as risk modeling studies seldom provide such information. As well, this specification is not directly aligned with the framework of Riley et al, which demands specifying outcome prevalence, c-statistic, and the calibration function. Fortunately, specification of our knowledge in terms of outcome prevalence, c-statistic, and calibration function, that is, defining $\theta=\{\phi,c,h\}$, can, under mild regularity conditions, fully identify $P(\pi,Y)$. The regularity conditions are as follows:
  1. 1)
    $h()$ is monotonically ascending (under the assumption of logit-linearity, this is satisfied as long as calibration slope is positive), and
  2. 2)
    $P(p)$ is quantile-identifiable; i.e., any two quantiles of the distribution are sufficient for uniquely identifying it. Typical distributions for risks, including Beta, Logit-normal, and Probit-normal satisfy this requirement23.
The first condition guarantees that the c-statistic relating $\pi$ to $Y$ (which is often reported) is equal to the c-statistic for $P(p)$ (as c-statistic is invariant under monotonical transformation of predictor values). The second condition is a requirement for the identifiability of $P(p)$ given its mean (prevalence) and c-statistic23.
As an example of this identifiability, consider an outcome prevalence of 0.25, c-statistic of 0.75. calibration slope of 1.1, and O/E ratio of 0.9. Assuming predicted risks have a Logit-normal distribution, the calibrated risks will also have Logit-normal distribution. The parameters of the latter are uniquely identifiable from $\phi,c$, which resolves to $P(p)\sim\mbox{Logitnorm}(-1.3302,1.0395)$ (the mcmapper R package implements our proposed numerical algorithms for this mapping24). As well, there is a 1:1 mapping between the O/E ratio and calibration intercept. Given the calibration slope of of 1.1 and the specified distribution for $p$, an O/E ratio of 0.9 uniquely maps to calibration intercept of -0.089 (see footnote of Table 1 for additional informaiton). Thus, to generate a random validation sample, one can sample $N$ calibrated risk from $P(p)$, generate corresponding response values as $Y_{i}\sim\mbox{Bernoulli}(p_{i})$ and compute predicted risks as $\pi_{i}=h^{-1}(p_{i})$. $(\pi_{i},Y_{i})$s created this way will be a realization of $D_{N}$.
Given this identifiability, characterizing our prior information involves specifying the joint distribution $P(\phi,c,h)$. Ideally, the previous analysis would report both the value and an estimated covariance matrix variance of $\hat{\theta}$, i.e., there would be joint inference about the three elements of theta. These would then become the prior mean and covariance matrix of the joint prior for $\theta$. In typical practice, however, joint inference on such parameters is not typically reported, but probability distributions for individual components can readily be constructed from existing information. For example, for outcome prevalence, if from a previous study with sample size of $n$, $m$ individuals experience the outcome, our knowledge can be specified as $\mbox{Beta}(m,n-m)$. For other components, point estimates and reported CI widths can be used, along an assumed distribution type, to construct distributions. Examples include specifying a Log-normal distribution for O/E ratio, Normal distribution for calibration slope, and Beta for c-statistic based on reported point estimates and bounds of 95%CI. The accompanying software provides a flexible way of specifying such distributions, accepting both direct specification of distribution parameters, moments, or mean and upper bound of 95%CI.
Once marginal distributions are constructed, a simple strategy would be to complete the prior specification by imposing a priori independence between the three components. An optional strategy, which would be more faithful to the true data generating mechanism, would involve a parametric bootstrap procedure to recover parameters interdependence. In this approach, one simulates multiple samples given $theta=\hat{thata}$, and for each sample record the ensuing estimates of $\hat{\theta}$. The empirical correlation matrix of these simulated estimates can then be taken as the prior correlation matrix, instead of simply presuming an uncorrelated prior. Appendix 1 provides an algorithmic description of this approach.
This algorithm in itself quantifies the uncertainty for a population that is exchangeable with the population(s) from which evidence on model performance is collected. If the evidence is collected from a single population, this specification assumes model performance is the same between the source and target populations. On the other hand, if current evidence is synthesized from multiple populations using meta-analytic techniques (as in our case study below), this specification assumes the target population is a random draw from the distribution of the meta-population. The predictive distribution of model performance in a new population can thus be used to characterize uncertainties. If the exchangability assumption does not hold, different steps of this algorithm can be modified to model population differences. If the outcome prevalence are expected to be different, one can shift the distribution of calibrated risks once they are determined in Step 4a to match the prevalence in the target population (which itself should be a random variable indicating our uncertainty about prevalence). Independently, if evidence is extracted from a model development study, and there are concerns about the model being overfitted, one can add a negative penalty term to the calibration slope in Step 4b (itself a random variable) representing our knowledge about the degree of overfitting. A structured examination of the degree of relatedness between the target population and source population(s) might help the investigator decide on the need for, and extent of, modifications25.
2.4.3 Sampling from posterior distributions
All sample size rules require expectation with respect to the distribution of parameters ($\theta$) and future sample ($D_{N}$). This naturally invites a Monte Carlo sampling algorithm based on repeated sampling from $P(\theta)$, simulating validation samples from $P(D_{N}|\theta)$, and computing the precision targets and VoI metrics in this sample. Draws from the posterior distributions are then processed according to the sample size rules of interest (e.g., average CI widths for the expected CI width criterion, the corresponding quantile for assurance-based metrics, or VoI for NB).
In addition to this default approach based on simulating $D_{N}$, we propose an approximate two-level closed-form approach for CI width targets that does not involve simulating $D_{N}$. This approaches is based on using SE equations twice. Let $\theta_{k}$ be the metric of interest (e.g., c-statistic) among $\theta$s. First, in the jth iteration of the Monte Carlo simulation, we specify $P(\hat{\theta}_{k}|\theta)$, the sample distribution of the metric of interest in the validation sample given our draw from $\theta$. This distribution is specified via method of moments with the first two moments being the corresponding value for $\theta_{k}$ in $\theta$, and $SE^{2}(\theta_{k})$ from the relevant SE equation, which is interpreted in a Bayesian flavor as the distribution of the future sample estimate. We then obtain a draw from such a distribution as a realization of $\hat{\theta}_{k}$, which is plugged into the SE equation again to quantify its CI width.
Table 1 provides an algorithmic description of both approaches.
Table 1: Bayesian Monte Carlo algorithm for drawing from the posterior distribution of precision targets
1. Assign a distribution to prevalence, c-statistic, and calibration function representing current knowledge (e.g., based on reported point estimates and confidence bands from previous development or validation studies). If calibration function is specified as a line on the logit scale, this can be one of the following. Distributions for calibration intercept and slope. Distributions for O/E ratio and calibration slope. Distributions for mean calibration and calibration slope. 2. Assign a distribution type for calibrated risks $p$ in the source population 3. Obtain a sample of size S for $\theta$: $\theta^{(j)}=\{\phi^{(j)},c^{(j)},h^{(j)}\},j=1,2,...,S$. Optional: use parametric bootstrapping to induce correlation among $\{\phi,c,h\}$ 4. For j=1 to S (number of Monte Carlo simulations). (a) Derive the parameters of $P^{(j)}(p)$, the distribution of calibrated risks in this iteration, given $\phi^{(j)}$ and $c^{(j)}$, given the distribution type assigned in step 2. *Sample-based method (b) Draw $N$ observations for calibrated risks: $p_{i}^{(j)}\sim P^{(j)}(p),i=1,2,...,N$. Draw corresponding response values $P(Y_{i}^{(j)})\sim\mbox{Bernoulli}(p_{i}^{(j)})$. Calculate the corresponding values of $\pi$: $\pi_{i}^{(j)}={h^{(j)}}^{-1}(p_{i}^{(j)})$. Construct ${D_{N}}^{(j)}=\{(\pi_{i}^{(j)},Y_{i}^{(j)})\}_{i=1}^{N}$ as the validation sample. (c) Using ${D_{N}}^{(j)}$, construct and record precision targets (CI widths): $ciw^{(j)}=g({D_{N}}^{(j)})$ for each metric of interest. *Two-step closed-form method (b) For any metric $\theta_{k}$, specify $P(\hat{\theta}_{k}|\theta^{(j)})$ using method of moments, with the first moment being true value of $\theta_{k}$ from $\theta^{(j)}$, and the second moment being $SE^{2}(\theta_{k})$ (and a choice of distribution type#). Obtain $\hat{\theta}_{k}$ as a draw from this distribution. (c) Plug $\hat{\theta}_{k}$ into the relevant SE equation to compute the precision criterion (CI width). Record this value. Next j 5. Process the draws from the precision targets according the sample size rule specified:
- For expected CI width criteria: $eciw=\sum_{j=1}^{S}ciw^{(j)}/S$.
- For quantile (assurance) CI width criteria: $qciw=\hat{F}_{ciw}^{-1}(q)$, where $\hat{F}_{ciw}$ is the empirical CDF of ciws and $q$ is the desired quantile (e.g., 0.9).
[…]
Currently, Beta, Logit-normal, and Probit-normal distributions are modeled in the accompanying R package.
#Central limit theorem justifies Normal distribution as the default. This is indeed compatible with the Wald method of constructing CI.
For VoI analysis, we are not aware of any previous work proposing optimality assurance. As well, the previously proposed algorithms for validation EVSI were fully Bayesian18. Our proposed algorithm for computing optimality assurance and EVSI for the mixed Bayesian-likelihood setup is presented in Table 2. We note that given that sample estimates of prevalence, sensitivity, and specificity all depend on the four frequencies $\{N_{tp},N_{fn},N_{tn},N_{fp}\}$ (the number of true positives, false negatives, true negatives, and false positives, respectively), $D_{N}$ can be fully specified by these four frequencies, which in turn can be drawn from their respective binomial distributions without simulating individual-level data.
Table 2: Computation of optimality assurance and EVSI
[…]
$N^{(j)}_{+}\sim\text{Binomial}(N,\phi^{(j)})$ (number of positive cases in the future sample),
$N^{(j)}_{tp}\sim\text{Binomial}(N^{(j)}_{+},se^{(j)})$,
$N^{(j)}_{fn}=N^{(j)}_{+}-N^{(j)}_{tp}$,
$N^{(j)}_{tn}\sim\text{Binomial}(N-N^{(j)}_{+},sp^{(j)})$,
$N^{(j)}_{fp}=N-N^{(j)}_{+}-N^{(j)}_{tn}$.
(e) Calculate sample esaimtes of NBs: ${NB}_{D_{N}^{(j)}}(k);k\in\{0,1,2\}$, by plugging in the sample estimates of prevalence, sensitivity, and specificity from ${D_{N}}^{(j)}$ in equation 2.2. (f) Identify the winning strategy in the sample: $W^{(j)}=\mbox{argmax}_{k}{NB}_{D_{N}^{(j)}}(k)$. (g) Record the true NB of this strategy $NBsample^{(j)}={NB}_{\theta}(W^{(j)})$. (h) Record whether the winning strategy had the highest possible NB: $A^{(j)}=I({NB}_{\theta}(W^{(j)})=NBmax^{(j)})$
Next $j$
6. Compute the proportion of times the winning strategy has the highest true NB: $Assurance=\sum_{j=1}^{S}A^{(j)}/S$. 7. Average NBs from step 5c: $ENB(k)=\sum_{j=1}^{S}NB^{(j)}(k)/S$. Pick the strategy that has the maximum ENB: $maxENB=\max_{k}ENB(k)$. This is the expected NB of the best strategy under current information. 8. Average $NBmaxs$ from step 5c: $ENBmax=\sum_{j=1}^{S}NBmax^{(j)}/S$. From this subtract $maxENB$. This is EVPI. 9. Average $NBsamples$ from step 5g: $ENBsample=\sum_{j=1}^{S}NBsample^{(j)}/S$. From this subtract maxENB. This is EVSI.
The integrals for sensitivity and specificity require numerical methods (with some exceptions, for example, for the Beta distribution).
2.4.4 Implementation
The above algorithms are implemented in the bayespmtools package (https://github.com/resplab/bayespmtools), in particular as two functions bpm_valprec() and bpm_valsamp(), for, respectively, computing precision / VoI for a fixed sample size, or determining the sample size corresponding to a set of precision / VoI criteria and sample size rules. As input, these functions expect evidence to be parameterized via four probability distributions, one for outcome prevalence, c-statistic, calibration slope, and one of calibration intercept, O/E ratio, or mean calibration. Missing correlation among these distributions is optionally (active by default) imputed via the parametric bootstrap method explained earlier. The correlation is induced to the sample of draws from marginal draws for these parameters via the methods by Iman and Conover26. This method offers the flexibility to the investigator to use different distribution types for each component (as opposed to a single multivariate distribution). Both algorithms currently implement Logit-normal, Probit-normal, or Beta for the distribution of calibrated risks.
bayespm_prec() computes expected CI widths, quantiles of CI widths, as well as optimality assurance and EVSI using the above-mentioned Monte Carlo sampling algorithms. bayespm_samp() solves for minimum $N$ that satisfies any of the requested precision criteria coupled with a sample size rule (e.g., targeting an expected CI width of 0.2 for O/E ratio, or 90% assurance that c-statistic CI width<0.1). Because for any $N$, one iteration of the Monte Carlo simulation generates one draw from the posterior distribution of CI widths or NBs, finding the minimum $N$ that satisfies a given sample size rule is a stochastic root-finding problem. In our implementation, for targets related to CI widths (expected value and assurance), we use the Robbins-Monro algorithm27, This was motivated by the simplicity of this algorithm which enables simultaneous optimization for all widths targets. For VoI and optimality assurance, we note that binomial draws can be vectorized across the entire sample of $\theta$s, thus these calculations can be performed faster than those based on CI Widths. We use the Stochastic-Simultaneous Optimistic Optimization algorithm by Valko et al28 (implemented in the OOR R package29).
3 Case study: Predicting deterioration in hospitalized COVID-19 patients
The ISARIC 4C model is a risk prediction model for predicting deterioration (need for ventilatory support or critical care, or death) in patients hospitalized due to COVID-19 infection. Details of the development and validation on this model are provided in its original publication30. In summary, the investigators used electronic hospital records from all 9 health regions of the UK to develop and validate . The London region was left out for external validation, and the remaining 8 regions were used in internal-external validation31. One at a time, one region was left out, the model was developed in the remaining 7 regions, and its performance was evaluated in the left-out region. Random-effects meta-analysis was used to pool out-of-sample estimates of discrimination and calibration metrics. Finally, the investigators used the data from all 8 regions to fit one final model, and evaluated its performance in the London region.
To use this setup as an informative example, we assume the London region did not participate in this study, and is now interested in conducting a validation study of this model in their region. That is, we ignore the results reported on the performance of the model in the London region, and consider the internal-external validation results as the information at hand. We assume that the performance of the model in the London region can be assumed to be a random draw from the distribution of performance observed across other regions (i.e., it is exchangeable with other regions).
The total development sample size was 70,349 (after removing those with unknown outcome status), with total number of events (deterioration) being 30,316, giving rise to an overall outcome prevalence of 0.428. Pooled estimates of the c-statistic, mean calibration, and calibration slope were 0.76 (95%CI 0.75–0.77), -0.01 (95%CI -0.12–0.09), and 0.99 (95%CI 0.97–1.02), respectively.
3.0.1 Application of Riley approach
As a baseline comparison, we performed sample size calculations for validating this model using the multi-criteria method proposed by Riley et al3. Riley et al used the ISARIC 4C as an example, and for consistency, we use the same targets: confidence interval widths of 0.1 for the c-statistic, 0.22 for O/E ratio, 0.3 for calibration slope. The assumed true values of these metrics was taken as the point estimates from the above-mentioned meta-analyses. The results indicate a sample size of 1,056 is required, which is dictated by the calibration slope. For other components, sample sizes are as follows: O/E ratio: 425, c-statistic: 359.
3.0.2 Application of Bayesian approach
Specifying $P(\theta)$: Given our assumption about the exchangeability of populations across UK geographic regions, the predictive distribution of the model performance in a new region represents our uncertainty about the model performance in the London region. However, it is crucial to note that pooled estimates and confidence bands reported in the original study are for the average effect and do not represent predictive distributions. To obtain predictive distributions, we re-performed the meta-analyses pooling c-statistic, mean calibration (note that while we specify evidence in terms of mean calibration, as is reported by Gupta et al, we target O/E ratio for sample size determination), and calibration slope for the 8 regions in the internal-external validation (we extracted study-specific estimates from digitized figures which enables us to use three significant digits). Gupta et al did not perform a meta-analysis on prevalence. We therefore performed this additional meta-analysis using region-specific prevalence data from their report. Given the large sample sizes, we performed all meta-analyses based on normality assumption on the original scale of each parameter, aside from c-statistic, for which a logit-transformation was performed as recommended by Snell et al32. We used method of moments to recover distribution from the mean and SD of the predictive distribution for a new population. We specified Beta distributions for prevalence and c-statistic, and Normal distributions for mean calibration and calibration slope. This resulted in the distributions for the four parameters presented in Table LABEL:tab:ISARIC_evidence.
Table 3: Evidence synthesis for the ISARIC study, based on predictive distribution of the internal-external validation results
Parameter Distribution Mean..SD. CI.bounds
prevalence Beta(119.64,159.91) 0.428 (0.03) 0.371–0.486
c-statistic Logit-normal(1.1565,0.0412) 0.761 (0.006) 0.746–0.775
mean calibration Normal(-0.0093,0.1245) -0.009 (0.125) -0.253–0.235
calibration slope Normal(0.9950,0.0237) 0.995 (0.024) 0.949–1.042
There are some approximations in the above approach that are made in favor of practicality. Possible correlations between parameters within each study were not available.We did not model between-study correlations (e.g., via a joint meta-analysis of parameters), after an exploratory analysis (examining Spearman rank correlation coefficients) did not support the presence of strong between-study correlation.
Target precision / VoI criteria and sample size rules: For comparability, we target the same CI width as in the Riley approach, targeting 95%CI width of, 0.22 for O/E ratio, 0.10 for c-statistic, and 0.30 for calibration slopes. For these metrics, we apply both the expected value and assurance rules. That is, we determine sample sizes that correspond to the expected CI width being equal to our targets (this can be interpreted as a direct Bayesian counterpart of conventional Riley et al’s formula). We also demand a 90% assurance to meet or exceed the CI width criteria, representing our preference for obtaining a narrower over wider CI width than the targeted with. As for clinical utility, we demand a 90% assurance on NB at the 0.2 threshold; that is, we desire to be 90% confident that the strategy that will emerge as the best in the validation sample is actually the best strategy in the target population. Finally, we investigate the expected gain in NB for each component of the sample size on the EVSI curve.
Analysis setup: All analyses are based on 10,000 draws from $P(\theta)$, assuming Logit-normal distribution for calibrated risks. This assumption is tested in a sensitivity analysis. We used the sample-based approach in this analysis. Results of the two-level closed-form approach are provided in Appendix 2.
Results: Table 4 provides the sample size for each component. The expected CIW components are close to their frequentist counterparts. The assurance-based targets result in slightly higher sample sizes, as expected. The final sample size of 1173 is dictated by the assurance component of the calibration slope.
Table 4: Sample size components based onconventional Riley approach (top row) and the Bayesian approach (bottom rows)
Approach c-statistic O/E ratio calibration slope NB
Conventional Riley approach (expected CI widths) 359 425 1056 149
Bayesian (expected CI width) 360 426 1025
Bayesian (90% assurance) 397 516 1173 306
Figure 1 shows how the precision targets change with sample size. These are computed from a separate Monte Carlo simulation of size 10,000 (independently of the calculations that determined the sample sizes) and as such act as diagnostic plots of CIW and NB assurance.
Figure 1: Expected CI widths (a) 90th quantile of CI width (b), optimality assurance (c) curves. Shapes on the line pertain to individual sample size components.
Refer to caption Refer to caption Refer to caption
(a) Expected value of CI width (b) 90th percentile of CI widths (c) Optimality assurance
Figure 2 shows the exemplary kernel histograms of the distribution of CI widths (the first three panels) for the smallest ($N=306)$ and largest ($N=1,173$) components of the sample size. The last panel demonstrates the distribution of the incremental NB of the model compared with the default strategies (i.e., $NB_{1}-\max(NB_{0}-NB_{2})$). With higher sample sizes, the CI widths get both shorter and also get more clustered. The distributions are relatively symmetrical. This can explain why the eciw values are close to their conventional, frequentist counterparts. For NB, as the sample estimate of NB is an unbiased estimator, higher sample sizes will result in a narrower distributions. The assurance for NB, and for EVSI, what affects the results is the share of the distribution that falls on the left side of zero.
Figure 2: Kernel histograms of CI widths for (a) c-statistic , (b) O/E ratio, and (c) calibration slope. Also, (d) shows the kernel histogram of the incremental net benefit of the model compared to the best defualt strategy for two sample sizes
Refer to caption Refer to caption
Refer to caption Refer to caption
Turning to VoI analysis, Figure 3 provides the EVSI calculation (with EVSI/EVPI ratio as the secondary Y-axis). Components of sample size calculation are overlaid on the graph.
Figure 3: Expected Value of Sample Information (EVSI) curve. Shapes pertain to individual sample size components
EVSI quantifies the expected value of a validation sample of a given size to identify the strategy with highest clinical utility
From this figure, one can (subjectively) realize that the CIW components fall on the ‘diminishing return’ of the EVSI. For example, going from the the largest sample size based on criteria other than the calibration slope (N=516) to the sample size dictated by assurance on calibration (N=1,173), a more than doubling of the sample size, is associated with 9.2% expected gain in NB. This can justify relaxing the criteria on calibration sloe for the validation study, if the focus is on clinical utility.
The decision whether to relax the calibration slope criteria can also be examined via the stability of the flexible calibration plots. As the desired precision based on the calibration slope is difficult to justify, but often dictates the final sample size required, Riley et al. recommend to plot the corresponding variability in calibration curves that might arise for that sample size33. In our context, the variability of a calibration curves has two sources: the variability in the true calibration function (due to variability in $\theta$) and the variability due to the finite validation sample (due to variability of $D_{N}$). The former is not a function of sample size. Because in our Bayesian Monte Carlo sampling algorithm, the true calibration function is known within each iteration, we can remove variability due to $\theta$ to quantify the difference between the flexible calibration curve and the true calibration function: for each simulated $D_{N}$, we fit a kernel regression (e.g., loess) for $P(Y=1|\pi)$, then remove $h(\pi)$ from the fitted values. Figure 4 shows such Bayesian calibration error plots for the final sample sizes including calibration slope criteria (N=1,173) and after removing such criteria (N=516). Moving to the smaller sample size expectedly increases the spread of calibration errors. Nevertheless, aside from extreme values of predicted risk, the 95% credible intervals (dashed lines) indicate that the error will mostly be <0.1. Again, this can be taken as a further argument in favor of settling on the smaller sample size.
Figure 4: Bayesian calibration error plots - dashed lines are 95% credible bands
Refer to caption Refer to caption
4 Discussion
We proposed a Bayesian framework for sample size considerations when a risk prediction model is to be evaluated (validated) in a target population. Compared to the conventional frequentist framework, this approach offers several advantages: it allows investigators to specify their uncertainties about model performance; it enables assurance-type sample size rules that incorporate investigators’ risk preferences; and it facilitates the the use of Value-of-Information analysis when clinical utility of the model is being examined. We proposed characterizing uncertainties around common metrics of model performance, and introduced Monte Carlo sampling algorithms for computing precision and VoI metrics.
Overall, this framework can provide an objective first step when designing validation studies aimed at quantifying the performance of a pre-specified model in a new population and determining whether it provides clinical utility. In this context, this framework helps the investigator explore the trade-off between sample size, precision, and gain in clinical utility. This framework can be used to investigate the anticipated precision or expected gain in clinical utility when the sample size is fixed (based on already collected data). Moroever, it can also be used to determine the sample size when original data collection is planned. A hybrid use, determining the sample size based on certain criteria and examining the resulting precision and VoI for other outcomes, is also possible. Our case study demonstrated such a hybrid approach. We defined rules targeting expected value and 90% assurance on the widths of confidence intervals around metrics of discrimination and calibration, as well as 90% assurance that we will be able to correctly identify the strategy with the highest clinical utility. Then, based on the results of the VoI analysis (as well as visual inspection of calibration errors), we could argue that one could potentially relax one component of the samples size requirement (precision around calibration slope), without losing much precision or clinical utility.
Among the outputs of this Bayesian approach, the assurance-type rules and VoI metrics are particularly relevant for risk prediction modeling. If the estimates of model performance are too uncertain, the adoption of the model might be met with resistance, even if the point estimates are compatible acceptable performance. This, combined with the general risk-averse attitude of public funding agencies34, means the costs of not meeting the desired targets is typically more than the benefits of exceeding those targets. Assurance-type rules can be used alongside those based on expected values to assuage concerns that the planned study might generate ambiguous results. Further, VoI metrics address longstanding criticisms on the irrelevance of inferential statistics when deciding on whether to adopt a health technology (including markers and risk scoring tools) for patient care9. These criticism have been recently voiced in the particular case of NB calculations for risk prediction models10. Accordingly, we recommend sample size considerations around NB to move away from confidence intervals and focus on assurance and EVSI. In particular, EVSI combines the risk of failing to identify the optimal strategy with the consequences (NB losses) of such a failure, providing a fuller picture of the implications of learning from an empirical study. We note that EVSI captures the expected gain in NB per decision. When scaled to the population (i.e., multiplied by the expected number of time the decision of interest is to be made), it quantifies the total value of conducting a validation study of a given size in net true or false positive units (an example of such calculations are provided in a previous study18). This provides a ‘value-based’ perspective that can complement the precision-based approach for sample size determination around metrics of discrimination and calibration.
There are several areas for further inquiry. We focused on binary outcomes, but this methodology can be extended to other outcome types. The application to survival outcomes seems particularly relevant and feasible. For such outcomes, model performance needs to be assessed at a time point of interest35. The time-dependent equivalents of prevalence and c-statistic are, respectively, the complement of survival probability and time-dependent area under the receiver operating characteristic curve at the time horizon of interest36. If uncertainty around these metrics are specified, our identifiability conditions can then be employed to map such metrics to the distribution of calibrated risks at this time point. However, the impact of censored observations in the future sample and the presence of competing risks on precision metrics and VoI values remains to be investigated. We focused on precision and VoI targets that pertain to the population-level performance of the model. Sample size consideration can also be approached in terms of uncertainty at the individual level, such as instability at individual-level predictions37. In line with Riley equations, we modeled the calibration function to be linear on the logit scale. This can be relaxed, for example by introducing a quadratic term as long as one can assure monotonicity of $h()$ and specify the joint distribution of the three parameters that would define it. A more appealing extension would be to model $h()$ non-parametrically. This can be done based a non-parametric summaries of calibration curves, such as those based on kernel smoothing (e.g., $ICI$16) or cumulative plots (e.g., $C^{*}$38).
From an implementation perspective, Bayesian calculations are inherently more complex than their frequentist counterparts. Several components of this framework require numerical integration and optimization methods. Examples include finding the parameters of $P(p)$ given prevalence and c-statistic, mapping O/E ratio or calibration mean to calibration intercept, and computing sensitivity and specificity (for NB calculations) at the threshold of interest given $P(p)$ and $h()$. These numerical algorithms might struggle for extreme cases. In addition, as the Bayesian Monte Carlo sampling produces draws from the posterior distributions of precision targets (e.g., CI widths), sample size determination requires stochastic optimization techniques, for example finding the $N$ that corresponds to the first moment or a quantile of this distribution. These algorithms demand convergence assessment and if needed repetition with different starting values. While our implementation performed robustly in our examinations (including the case study), numerical accuracy and stability of these algorithms should be considered across a range of input values in dedicated studies. We consider the accompanying software to be an evolving implementation, subject to revisions and improvements.
Compared to the conventional sample size determination methods, the Bayesian approach is predicated on one crucial extra step: characterizing our uncertainties about model performance. Such characterization is not currently a common practice in applied prediction modeling. The conventional wisdom in risk modeling studies is to report a ‘final model’ as a deterministic function that maps patient characteristics to an exact value of conditional outcome risk. This practice ignores the fact that predictions for a model trained using a finite sample are inherently uncertain. Differences across populations in the relationship between predictors and the outcome, differences in predictor assessment approaches across settings, and variations in outcome ascertainment methods further add to our uncertainties39. Fortunately, there have been recent calls on the importance of characterizing and communicating uncertainty around model predictions40. These calls encourage investigators to fully document and communicate uncertainty in model coefficients in reports of model development (e.g., reporting the covariance matrix of model coefficients, or reporting the coefficients of models fitted in bootstrapped copies of the original sample)40. These can further facilitate incorporating uncertainties into the design of subsequent studies.
A thought-provoking consequence of adopting a Bayesian approach is the ultimate fate of the prior information. Reflecting current practices, we assumed the future validation sample will be the sole source of evidence on model performance once it is procured, in effect discarding our current knowledge once it is used to construct $P(\theta)$. But, if our existing knowledge is informative enough to warrant investigating the model in a new setting, should it not be used alongside the sample to inform our final judgment? Taking our case study as an example, the information on prevalence alone (based on the moments of its distribution) is equal to having learned its value from a sample of 280 individuals - a non-trivial amount of information! Incorporating such prior knowledge into what we learn from the data will reduce prediction uncertainty, which can affect both patient care (more precise predictions) and the design of empirical studies (requiring smaller samples to achieve the same targets). Of course, a perceived drawback is the subjectivity inherent in real-world Bayesian reasoning. Our case study might be seen as a stylized setup where predictive distributions from a meta-analysis were available and the assumption of exchangability of populations seemed tenable. Things can be more subjective, for example, when a single development study in another population is the sole source of evidence. Our current response to the fear of compromised objectivity is to completely drop the existing information. Perhaps one can instead define objectivity in terms of explicit and transparent specification of prior evidence and steps taken to account for population heterogeneity.
Overall, a cultural shift towards embracing uncertainty in predictions may also encourage the adoption of study designs that more effectively leverage existing knowledge. We expect Bayesian approaches towards study design to gain traction as awareness grows regarding the relevance of prediction uncertainty in decision making.
5 Appendix
5.1 Appendix 1: Algorithm for parametric bootstrapping
1. For i in 1 to M (size of the Monte Carlo simulation) (a) Generate $D_{n}:=\{(\pi_{i},Y_{i})\}_{i=1}^{n}$, an iid sample of size $n^{\dagger}$ consisting of predicted risks and response values using the point estimates $\hat{\theta}=\{\hat{\phi},\hat{c},\hat{h}\}^{\ddagger}$. (b) Estimate $\theta^{(j)}=\{\phi^{(j)},c^{(j)},h^{(j)}\}$ from the above sample. 2. Estimate the correlation coefficients for $\theta$s. Generally, correlation coefficients are not sensitive to the size of the sample. It is natural to choose $n$ to be the size of the original sample. In instances (as in our case study) where elements of $\theta$ are taken from different sources or from the output of meta-analysis (as in our case study), one can determine $n$ based on the effective sample size (e.g., based on uncertainty around prevalence). DEtermining the distribution of predicted risks and observed responses from .
5.2 Appendix 2: Sample size and CI widths for the two-level closed-form method
Approach c-statistic O/E ratio calibration slope
Bayesian (expected CI width) 343 420 1044
Bayesian (90% assurance) 398 516 1173
Expected CI widths (a) 90th quantile of CI width (b), optimality assurance (c) curves based on the two-level closed-form method. Shapes on the line pertain to individual sample size components.
Refer to caption Refer to caption
(a) Expected value of CI width (b) 90th percentile of CI widths

### v2

@@
 Abstract
-Background: Contemporary sample size calculations for external validation of risk prediction models require users to specify fixed (i.e., true) values of assumed model performance metrics alongside target precision levels (e.g., 95% CI widths).
+Contemporary sample size calculations for external validation of risk prediction models require users to specify fixed values of assumed model performance metrics alongside target precision levels (e.g., 95% CI widths).
 However, due to the finite samples of previous studies, our knowledge of true model performance in the target population is uncertain, and so choosing fixed values represents an incomplete picture.
 As well, for net benefit (NB) as a measure of clinical utility, the relevance of conventional precision-based inference is doubtful.
-Bayesian approach can address this, by accounting for uncertainty of true model performance and facilitating decision-theoretic metrics including the expected value of sample information (EVSI).
-Methods: We propose a general Bayesian algorithm for constructing the joint distribution of predicted risks and response values based on summary statistics of model performance in previous studies.
-For statistical metrics of performance, we propose sample size determination rules that either target desired expected precision, or a desired assurance probability that the precision criteria will be satisfied.
-For NB, we propose rules based on optimality assurance (the probability that the planned study correctly identifies the most beneficial strategy) and the EVSI, the expected gain in net benefit from the planned validation study.
+In this work, we propose a general Bayesian framework for multi-criteria sample size considerations for prediction models for binary outcomes.
+For statistical metrics of performance (e.g., discrimination and calibration), we propose sample size rules that target desired expected precision or desired assurance probability that the precision criteria will be satisfied.
+For NB, we propose rules based on Optimality Assurance (the probability that the planned study correctly identifies the optimal strategy) and Value of Information (VoI) analysis.
 We showcase these developments in a case study on the validation of a risk prediction model for deterioration of hospitalized COVID-19 patients.
-Results: The contemporary approach, based on fixed values of assumed c-statistic, O/E ratio, and calibration slope from in the validation study, with target 95%CI width of, respectively, 0.10, 0.22, and 0.30, would recommend a sample size of 1,056, dictated by the desired precision around the calibration slope.
-To implement the Bayesian approach and to account for the uncertainty of the model’s predictive performance, we constructed predictive distributions for prevalence, c-statistic, mean calibration error, and calibration slope for a new validation study.
-Targeting the same expected CI width would result in a sample size of 1,025.
-However, demanding 90% assurance for meeting these precision criteria would result in a sample size of 1,173.
-In terms of NB, an optimality assurance of 90% was achieved at a sample size of 306.
-The EVSI curve demonstrated a diminishing margin beyond sample sizes >500 (<10% gain in expected NB with doubling the sample size).
-Conclusion: Compared to the conventional sample size calculation methods, a Bayesian approach requires explicit quantification of uncertainty around model performance, but thereby enables various sample size rules based on expected precision, assurance probabilities, and value of information.
-In our case study, EVSI calculations indicated that the precision criterion around calibration slope could potentially be relaxed without much utility loss.
-1 Background
+Compared to the conventional sample size calculation methods, a Bayesian approach requires explicit quantification of uncertainty around model performance, and thereby enables flexible sample size rules based on expected precision, assurance probabilities, and VoI.
+In our case study, calculations based on VoI for NB suggest considerably lower sample sizes are needed than when focusing on precision of calibration metrics.
+footnotetext: From Faculty of Medicine and Faculty of Pharmaceutical Sciences, The University of British Columbia (MS); Department of Statistics, The University of British Columbia (PG); British Columbia Centre For Disease Control (SS); Department of Epidemiology, CAPHRI Care and Public Health Research Institute, Maastricht University, and Department of Development and Regeneration, KU Leuven (LW); Department of Applied Health Sciences, School of Health Sciences, College of Medicine and Health, University of Birmingham, and National Institute for Health and Care Research (NIHR) Birmingham Biomedical Research Centre (RDR)footnotetext: $\dagger$ Co-senior authors with equal contributionfootnotetext: * Correspondence to Mohsen Sadatsafavi, Room 4110, Faculty of Pharmaceutical Sciences, 2405 Wesbrook Mall, Vancouver, BC, V6T1Z3, Canada; email: msafavi@mail.ubc.ca
+Introduction
 Developing risk prediction models or validating existing ones in a new population represents significant investment in time, resources, and expertise.
@@
 A particular aspect of study design is the sample size of such studies.
-The field has witnessed significant recent developments on this front1–8.
-An example is the multi-criteria approach by Riley et al for the sample size required for external validation studies, targeting, targeting pre-specified width of the 95% confidence interval around metrics of model performance2,3.
+The field has witnessed significant recent developments on this front[1, 2, 3, 4, 5, 6, 7, 8].
+An example is the multi-criteria approach by Riley et al for the sample size required for external validation studies, targeting pre-specified widths of the 95% confidence intervals around metrics of model performance[2, 3, 9].
 For binary outcomes, such metrics are related to discrimination, calibration, and net benefit (NB).
-The sample size required for each component is computed separately, with the largest one advertised as the final requirement3.
-In addition to target precision criteria, each component of this approach requires as input an assumed true value of the the metric of interest in the target population.
-In reality, we do not know the true value of model performance metrics with certainty (otherwise there would be no need to conduct the study in the first place).
+The sample size required for each component is computed separately, with the largest one advertised as the final requirement[3].
+In addition to target precision criteria, each component of this approach requires as input an assumed true value of the metric of interest in the target population.
+In reality, we do not know the true value of model performance metrics with certainty (otherwise there would be no need to conduct a validation study in the first place).
 Further, due to sampling variability, the precision obtained in one particular dataset may differ substantially from the expected precision for that sample size.
 Thus, there is more uncertainty to be accounted for than the Riley criteria allow.
-A Bayesian approach towards sample size determination allows uncertainty to be accounted for, which is advantageous on multiple fronts.
+A Bayesian approach towards sample size determination allows this uncertainty to be accounted for, which is advantageous on multiple fronts.
 First, it enables the full use of existing information on model performance at the time the study is designed, rather than forcing the investigator to express their knowledge as the fixed truth.
-Second, it enables different classes of sample size rules, namely those that target the expected values of precision targets, as well as ‘assurance’-type rules for the probability of meeting (or exceeding) precision targets.
-This can be insightful as focusing on expected values alone does not guarantee that the desired precision will be achieved in one particular dataset, and the investigator may want a stronger assurance against this.
-For example, an investigator that desires an interval width of 0.1 for c-statistic might not perceive much benefit from the future interval being narrower, but might have a strong preference against not meeting this target.
-As such, targeting a sample size that will result in 90% probability that the CI width will be $\leq$ 0.1 might be more appealing than a sample size that targets an expected CI width of 0.1.
-Finally, when assessing clinical utility, the relevance of precision-based criteria is challenged9,10.
-A Bayesian approach enables the use of novel, decision-theoretic approaches based on value of information (VoI) analysis, which focus on the expected gain in clinical utility from increasing sample size11.
-Bayesian approaches for power and sample size calculations are mostly investigated in the context of experimental studies that are aimed at interrogating a null hypothesis12–15.
+Second, it enables users to choose from different classes of sample size rules, namely those that target the expected values of precision targets (as in Riley criteria), as well as ‘assurance’-type rules for the probability of meeting (or exceeding) precision targets.
+The latter can be insightful as focusing on expected values alone does not guarantee that the desired precision will be achieved in one particular dataset, and the investigator may want a stronger assurance against this.
+For example, an investigator that desires an interval width of 0.1 for the c-statistic might not perceive much benefit from the future interval being narrower, but might have a strong preference against not meeting this target.
+As such, targeting a sample size that will result in, say, a 90% probability that the CI width will be $\leq$ 0.1 might be more appealing than a sample size that targets an expected CI width of 0.1.
+Finally, when assessing clinical utility, the relevance of precision-based criteria is challenged[10, 11].
+A Bayesian approach enables the use of novel, decision-theoretic approaches based on Value of Information (VoI) analysis, which focuses on the expected gain in clinical utility from increasing sample size[12].
+Bayesian approaches for power and sample size calculations are mostly investigated in the context of experimental studies that are aimed at interrogating a null hypothesis[13, 14, 15, 16].
 However, risk model validation studies are not generally hypothesis-driven.
 Rather, the focus is on the precise estimation of a variety of metrics of model performance.
-As such, Bayesian sample size considerations in this context is worth exploring.
-Hence, in this article, we propose a Bayesian version of the sample size formula by Riley et al, focusing on the same metrics of model performance, but enabling the investigator to 1) incorporate their uncertainty around assumed model performance in calculations, 2) use sample size rules that target assurance probabilities for meeting chosen criteria, and 3) use rules based on VoI analysis for clinical utility.
+As such, Bayesian sample size considerations in this context are worth exploring.
+Hence, in this article, we propose a Bayesian version of the sample size formula by Riley et al, which enables the investigator to 1) incorporate their uncertainty around assumed model performance in calculations, 2) use sample size rules that target assurance probabilities for meeting chosen criteria, and 3) use rules based on VoI analysis for clinical utility.
 The proposed framework can be used in two general ways: to quantify the anticipated precision or VoI outcomes from a planned study if the sample size is fixed (as is the case with validation studies that are based on already collected data), or to determine the minimum sample size that achieves pre-specified precision or VoI criteria.
 This framework can also be used in a hybrid form: determine the sample size based on certain criteria (e.g., targeting CI width for c-statistic and calibration slope), and investigate the consequence of the chosen sample size using other criteria (e.g, assurance probability around NB for the final sample size).
-Calculations can be done for the entire sample, as well as within subgroups, for example, imposing fairness criteria around the precision of estimates among minority sub-groups.
+Calculations can be done for the entire sample, as well as within subgroups, for example, imposing fairness criteria around the target precision of estimates among minority sub-groups.
 The rest of this manuscript is structured as follows.
-First, we briefly review common metrics of model performance and multi-criteria sample size formula by Riley et al.
-We will then introduce a Bayesian extension of this framework based on various sample size rules.
-We propose a general approach for characterizing uncertainty around model performance based on commonly reported information from previous studies.
-We outline Monte Carlo sampling algorithms for drawing from the posterior distributions of relevant quantities.
+First, we briefly review common metrics of model performance and multi-criteria sample size formula by Riley et al[3].
+We will then introduce our Bayesian extension of this framework based on various sample size rules.
+We propose a general approach for characterizing uncertainty around model performance based on commonly reported information from previous studies, and outline Monte Carlo sampling algorithms for drawing from the corresponding distributions.
 A case study based on a model for predicting COVID-19 deterioration showcases the developments.
 We conclude by suggesting further areas for research.
-2 Methods
-We review the context, common metrics of model performance, Riley’s sample size formulas for external validation studies, and our proposal for Bayesian sample size determination.
-2.1 Context
+Methods
+In this section we review the context, common metrics of model performance, Riley’s sample size formulas for external validation studies, and our proposal for Bayesian sample size determination.
+Context
 We focus on external validation of a model for predicting the risk of a binary outcome.
 Broadly speaking, an external validation study is an endeavor in learning about the joint distribution of predicted risks ($\pi$) from a pre-specified model and observed outcomes ($Y$: 0 no-event, 1 event) in a target population.
-The validation sample $D_{N}$ of size $N$ can be seen as $N$ pairs of predicted risks and observed results: $D_{N}=\{(\pi_{i},Y_{i})\}_{i=1}^{N}$.
-A classical validation study does not involve learning about the relationship between predictors and the outcome.
-Rather, the focus is on quantifying the performance of a pre-specified model, and ultimately identifying whether it is worth implementing in the target population.
-For ease of expositions, we assume one model is being considered.
+The validation sample $D_{N}$ of size $N$ participants can be seen as $N$ pairs of predicted risks and observed results: $D_{N}=\{(\pi_{i},Y_{i})\}_{i=1}^{N}$.
+A validation study does not involve learning about the relationship between predictors and the outcome.
+Rather, the focus is on quantifying the performance of an existing model.
+For ease of expositions, we assume one model is being evaluated.
 This framework can be extended to multi-model comparisons with relative ease, but we leave this to subsequent exploration.
 As is the standard in contemporary practice, predicted risks for the model are taken as fixed values, representing the expected outcome probability among all individuals with the same predictor pattern.
-The proposed approach is applicable for both classical regression models and black-box machine learning algorithms, as a validation study is not concerned on how predictors are used to compute the predicted risk.
-2.2 Common metrics of model performance
+The proposed approach is applicable whenever a model or algorithm provides estimated risks for individuals (regardless of whether it was developed using regression, machine learning or AI-based methods), as a validation study is not concerned on how predictors are used to compute the predicted risk.
+Common metrics of model performance
 Most model development and validation studies report the following metrics along with associated uncertainty:
@@
 This function converts the predicted risk for the ith individual, $\pi_{i}$, to the true risk for that person, which we denote by $p_{i}$.
-$p_{i}:=h(\pi_{i})$.
-When using non-functional form $h$ we refer to the parameters that define $h()$.
+$p_{i}:=h(\pi_{i})$, which is the expected values of observed outcome values across all individuals with similar predictor patterns.
+As needed, by $h$ we refer to parameters describing $h()$.
 Commonly, including in Riley’s sample size formula, $h()$ is modeled linearly on the logit scale: $\mbox{logit}(h(\pi))=\alpha+\beta\mbox{logit}(\pi)$, and thus $h$ consists of an intercept ($\alpha$) and slope ($\beta$).
 Sometimes studies report $\beta$ and ‘mean calibration’, i.e., $\mathbb{E}(Y-\pi)$, or $\beta$ and observed-to-expected outcome ratio ($O/E:=\mathbb{E}(Y)/\mathbb{E}(\pi)$), but these are equivalent as for a fixed value of $\beta$, knowing any of calibration intercept, mean calibration, or O/E ratio is enough to specify the calibration line.
-In other instances, $h()$ is modeled more flexibly using non-parametric methods (e.g., based on LOESS smoothing)16.
-Riley et al, while determining sample size based on a logit-linear calibration function, suggests visually inspecting the variability in anticipated smoothed curves once the sample size is determined2.
+In other instances, $h()$ is modeled more flexibly using non-parametric methods (e.g., based on LOESS smoothing)[17].
+Riley et al, while determining sample size based on a logit-linear calibration function, suggests visually inspecting the variability in anticipated smoothed curves once the sample size is determined[2].
 3) c-statistic $(c)$: This is a measure of the discriminatory performance of the model, and is the probability of concordance between the ranking of predicted risks and outcomes among a randomly chosen pair.
 Formally, $c:=P(\pi_{2}>\pi_{1}|Y_{2}=1,Y_{1}=0)$ with $(\pi_{1},Y_{1})$ and $(\pi_{2},Y_{2})$ being two randomly selected pairs of predicted risks and responses.
-4) Net benefit (NB): Details of NB calculations are provided elsewhere17.
+4) Net benefit (NB): Details of NB calculations are provided elsewhere[18].
 In a nutshell, at a chosen risk threshold $z$, there are at least three treatment strategies concerning the use of the model: treat no one (with a default NB of 0), use the model to treat individuals whose predicted risk is $\pi\geq z$, or treat all.
@@
 On the contrary, metrics of model performance such as c-statistic or calibration function do not consider the decision-making context.
-2.3 Riley’s sample size formulas
-Citing the commonality of the above-mentioned metrics in risk modeling studies, Riley et al structured their multi-criteria equations around desired precision levels (width of 95%CI) around these metrics3.
+Riley’s sample size formulas
+Citing the commonality of the above-mentioned metrics in risk modeling studies, Riley et al structured their multi-criteria equations around desired precision levels (width of 95%CI) around these metrics[3].
 In particular, the following equations were suggested for approximating the standard deviation of the sampling distribution of the estimator (i.e., standard error [SE]), from which Wald-type confidence intervals could be constructed:
@@
 where $I_{\alpha}={\mathbb{E}}\left(p(1-p)\right)$, $I_{\beta}={\mathbb{E}}\left(\mbox{logit}(\pi)^{2}p(1-p)\right)$, and $I_{\alpha,\beta}={\mathbb{E}}\left(\mbox{logit}(\pi)p(1-p)\right)$, and $p=h(\pi)$.
-Assuming all measures are of interest, the final sample size is decided by the largest $N$ among each component.
+Assuming all measures are of interest, the final sample size is decided by the largest $N$ among each component (and by checking whether the subsequent variability of calibration curves is acceptable - as explained above).
 Riley et al proposed also targeting desired CI bands around standardized NB.
-However, a Bayesian approach facilitates using decision-theoretic approaches for NB, addressing the criticisms around the relevance of conventional inferential methods for NB9,10.
+However, a Bayesian approach facilitates using decision-theoretic approaches for NB, addressing the criticisms around the relevance of conventional inferential methods for NB[10, 11].
 Note that the calculations require assumptions about the anticipated model performance (e.g., c-statistic, calibration slope) and the distribution of predicted risks in the target population.
@@
 We now propose a Bayesian approach that enables modeling uncertainties around these quantities.
-2.4 Going Bayesian
-A Bayesian approach towards sample size calculation requires 1) identifying metrics of interest for model performance that will be estimated from this sample (e.g., c-statistic, calibration slope), 2) deciding on precision targets for such metrics (e.g., a 95% CI width of 0.1 for c-statistic), 3) deciding on sample size rules on such metrics (e.g., targeting the expected CI width, demanding a 90% assurance that the CI width will be met, or 90% assurance that the future validation study will correctly identify the most optimal decision at a risk threshold of interest; 4) specifying our current information about model performance; 5) generating random draws from the posterior distribution of desired targets that will be processed according to the specified rules.
+Going Bayesian
+A Bayesian approach towards sample size calculation requires 1) identifying metrics of interest for model performance that will be estimated from this sample (e.g., c-statistic, calibration slope); 2) deciding on precision targets for such metrics (e.g., a 95% CI width of 0.1 for c-statistic); 3) deciding on sample size rules on such metrics (e.g., targeting the expected CI width, demanding a 90% assurance that the CI width will not be larger than targeted, or 90% assurance that the future validation study will correctly identify the most optimal decision at a risk threshold of interest); 4) specifying our current information about model performance; 5) generating random draws from the distribution of desired targets that will be processed according to the specified rules.
 In line with the contemporary practice in risk prediction development and validation, we adopt a mixed Bayesian-likelihood perspective: we use existing evidence (prior) to quantify our uncertainty about model performance, but we assume once the future sample is obtained, we will solely rely on it (the likelihood).
@@
 This joint distribution is indexed by parameters $\theta$, a random entity summarizing our current knowledge about this joint distribution (which encompasses our beliefs about model performance in the target population).
-We learn about $\theta$ by collecting ${D}_{N}$.
+We plan to learn about $\theta$ by collecting a sample of data $D_{N}$ from $N$ participants in the target population.
 The key aspect of a Bayesian approach is that we treat $\theta$, $D_{N}$, and therefore any quantity (e.g., CI widths) derived from $D_{N}$, as random entities.
@@
 Repeatedly, we sample from $P(\theta)$ and simulate a random validation sample $P(D_{N}|\theta)$, from which performance metrics are quantified and recorded.
-Repeating this many times will generate draws from the posterior distribution of these performance metrics for a validation dataset of a particular sample size.
-These can be summarized in terms of their mean (expected value) and variance, and lead to various sample size rules and assurance probabilities.
-Naturally, Bayesian Monte Carlo sampling results in draws from posterior distributions of targets (e.g., CI widths) given a fixed $N$.
+Repeating this many times will generate draws from the prior predictive distributions (henceforth referred to as ‘pre-posterior’ distribution, the anticipated distribution of the quantity of interest after the future sample is procured) of these performance metrics for a validation dataset of a particular sample size.
+These draws can be processed according to various sample size rules and assurance probabilities.
+Naturally, Bayesian Monte Carlo sampling results in draws from pre-posterior distributions of targets (e.g., CI widths) given a fixed $N$.
 Sample size calculation thus becomes a stochastic inverse problem: find the smallest $N$ that satisfies a set of sample size rules.
-As the latter is detached from the core of a Bayesian workflow, we primarily focus on calculating precision and VoI metrics for a given $N$, and discuss our implementation of stochastic root-finding algorithms for sample size determination afterwards.
-2.4.1 Sample size rules
+As the latter is detached from the core of a Bayesian workflow, we primarily focus on calculating precision and VoI metrics for a given $N$, but also briefly discuss our implementation of stochastic root-finding algorithms for sample size determination.
+Sample size rules
 Because the Bayesian approach takes estimates of model performance in the future validation study as random quantities, it invites establishing sample size rules that take into account such randomness.
@@
 For NB as a measure of clinical utility, we suggest sample size rules that are not precision-based and rather hone in on our ability in detecting the strategy with the highest NB.
-2.4.1.1 Sample size rules for metrics of discrimination and calibration
+Sample size rules for metrics of discrimination and calibration
 Broadly speaking, we consider some scalar summary derived by applying summarizing function $g()$ to validation data ${D}_{N}$, to meet some criterion.
 In our context, $g({D}_{N})$ returns the 95%CI width for the corresponding metric.
-Expected CI widths (eciw): This is related to the Average Length Criterion discussed by Joseph et al13. Here, we target the expected CI width (eciw) across the distribution of ${D}_{N}$:
-$eciw(N)=\mathbb{E}_{\mathbf{\theta}}\mathbb{E}_{{D}_{N}|\mathbf{\theta}}(g({D}_{N}))$
+Expected CI widths (ECIW): This is related to the Average Length Criterion discussed by Joseph et al[14]. Here, we target the expected CI width (ECIW) across the distribution of ${D}_{N}$:
+$ECIW(N)=\mathbb{E}_{\mathbf{\theta}}\mathbb{E}_{{D}_{N}|\mathbf{\theta}}(g({D}_{N})).$
 If the sample size is fixed, $ECIW(N)$ is reported for each component as the expected precision of the future study.
-For sample size calculation, we seek the minimum $N$ that results in $eciw(N)$, meeting precision target $\tau$: $\min\{N\in\mathbb{N}|eciw(N)\leq\tau\}$.
+For sample size calculation, we seek the minimum $N$ that results in $ECIW(N)$ meeting precision target $\tau$: $\min\{N\in\mathbb{N}|ECIW(N)\leq\tau\}$.
 We derive $N$s separately for each component, and choose the maximum $N$ as the final sample size.
-For example, we might identify the $N$ required to ensure that the expected CI widths of the calibration slope, and c-statistic all meet a desired expected CI width.
-Assurance-type rules based on quantiles of CI width (qciw): This is related to the modified worst outcome criterion as discussed by Joseph et al13.
+For example, we might identify the $N$ required to ensure that the expected CI widths of the calibration slope and c-statistic all meet a desired expected CI width.
+Assurance-type rules based on quantiles of CI width (QCIW): This is related to the modified Worst Outcome Criterion as discussed by Joseph et al[14].
 Here, we target the probability of meeting (or exceeding) desired precision targets.
@@
 with
-$F_{N}(q)=P(g(D_{N})\leq q)=\mathbb{E}_{\mathbf{\theta}}\mathbb{E}_{{D}_{N}|\mathbf{\theta}}(I(g({D}_{N})\leq q))$
+$F_{N}(x)=P(g(D_{N})\leq x)=\mathbb{E}_{\mathbf{\theta}}\mathbb{E}_{{D}_{N}|\mathbf{\theta}}(I(g({D}_{N})\leq x))$
 being the CDF of the distribution of the anticipated CI width for sample size $N$.
 Again, for a fixed-$N$ setup we report $QCIW(N)$.
-For sample size determination, we find the minimum $N$ such the the $q^{th}$ quantile is not greater than the target CI width $\tau$: $\min\{N\in\mathbb{N}|qciw(N,q)\leq\tau\}$.
-2.4.1.2 Sample size rules for NB
-Optimality assurance for NB : This is the expected probability that we will correctly identify the strategy that has the highest population NB based on $D_{N}$.
+For sample size calculation, we find the minimum $N$ such the the $q^{th}$ quantile is not greater than the target CI width $\tau$: $\min\{N\in\mathbb{N}|QCIW(N,q)\leq\tau\}$.
+Sample size rules for NB
+Optimality Assurance for NB : This is the probability that we will correctly identify the strategy that has the highest population NB based on $D_{N}$.
 To proceed, let $NB_{D_{N}}(k)$ be the sample estimate of $NB(k)$, and $NB_{\theta}(k)$ its true value given $\theta$.
 We note that with the future data at hand, the investigator will declare a winning strategy as the one that has the highest expected NB solely based on the sample:
-$W(D_{N})=\mbox{arg}\max_{k}(NB_{D_{N}}(k))$
+$W(D_{N})=\mbox{arg}\max_{k}(NB_{D_{N}}(k)).$
 Optimality Assurance is the probability that the NB of the strategy that we will declare as the best in the sample is the maximum possible NB:
 $Assurance(N)=P\left(NB_{\theta}(W(D_{N}))=\max_{k}(NB_{\theta}(k))\right)=\mathbb{E}_{\theta}\mathbb{E}_{D_{N}|\theta}I\left(NB_{\theta}(W(D_{N}))=\max_{k}(NB_{\theta}(k))\right).$
-This assurance is in the probability scale that is non-decreasing and asymptotic to 1 as the sample size is increased.
-Expected Value of Sample Information: The key VoI quantity is EVSI(N), the expected gain in NB from conducting a future validation study of size $N$.
+This assurance is in the probability scale that is non-decreasing and asymptotes to 1 as the sample size is increased.
+Expected Value of Sample Information (EVSI): The key VoI quantity is $EVSI(N)$, the expected gain in NB from conducting a future validation study of size $N$, compared with the NB of the decision made with current information:
 $EVSI(N)=\mathbb{E}_{\theta}\mathbb{E}_{D_{N}|\theta}\left(NB_{\theta}(W(D_{N}))\right)-\max_{{}_{k}}{\mathbb{E}_{\mathbf{\theta}}NB_{\theta}(k)}.$
-The first term on the right hand side is the expcted NB of the decision that we will declare as optimal based on $D_{N}$.
+The first term on the right-hand side is the expcted NB of the decision that we will declare as optimal based on $D_{N}$.
 The second term on the right-hand side is the expected NB of the best decision under current information.
@@
 $EVSI(N)$ is a non-decreasing function.
-Its maximum value occurs when $N$ is infinity, that is, we have access to the entire population and can unequivocally determine the best strategy.
-The expected gain from such perfect information is called the expected value of perfect information (EVPI) and can be calculated as (18,19):
+Its maximum value occurs when $N$ is infinity; that is, we have access to the entire population and can unequivocally determine the optimal strategy.
+The expected gain from such perfect information is called the Expected Value of Perfect Information ($EVPI$) and can be calculated as[19, 20]:
 $EVPI=\mathbb{E}_{\mathbf{\theta}}\max_{k}(NB_{\theta}(k))-\max_{k}{\mathbb{E}_{\mathbf{\theta}}NB_{\theta}(k)}.$
@@
 An $rEVSI$ of 0.4 for a given sample size means that the external validation study at this sample size is expected to reduce the expected NB loss due to uncertainty by 40%.
-Another sample size rule can be based on the expected net benefit of sampling (ENBS)20.
+Another sample size rule can be based on the Expected Net Benefit of Sampling ($ENBS$)[21].
 $ENBS$ requires scaling the EVSI by the expected size of the population affected by the decision, and subtracting the costs of collecting a sample in the same (true or false positive) unit as NB.
 Let $\mathcal{M}$ be the expected number of times the decision of interest is to be made, and $\mathcal{W}$ be the effort of every additional recruitment for the validation study in NB units.
-Then
+Then:
 $ENBS(N)=\mathcal{M}EVSI(N)-\mathcal{W}N,$
@@
 This rule for sample size determination is fully decision-theoretic as it avoids specifying any arbitrary threshold values.
-However, it requires scaling total population size and establishing the trade-off between sampling efforts and clinical utility; both of these tasks are context-specific.
+However, it requires scaling NBs to the population and establishing the trade-off between sampling efforts and clinical utility; both of these tasks are context-specific.
 As such, while we propose this rule here for completeness, we will not investigate it in the case study.
-2.4.2 Specifying $P(\theta)$
+Specifying $P(\theta)$
 The Bayesian approach requires specifying $P(\theta)$, the distribution of parameters that govern $P_{\theta}(\pi,Y)$.
@@
 There weights are random with a distribution of $P(\theta)\sim\mbox{Dirichlet}(1,...,1)$.
-The empirical distribution of this weighted sample can be considered as a random draw from the distribution of the population21.
+The empirical distribution of this weighted sample can be considered as a random draw from the distribution of the population[22].
 The validation sample $D_{N}$ then can be obtained from sampling with replacement.
-Details of this two-level resampling approach is provided elsewhere18,22.
+Details of such two-level resampling approach is provided elsewhere[23, 19].
 Our focus here is on parametric modeling based on summary statistics from previous studies.
@@
 However, specifying the distribution of predicted risks in the target population is not straightforward, as risk modeling studies seldom provide such information.
-As well, this specification is not directly aligned with the framework of Riley et al, which demands specifying outcome prevalence, c-statistic, and the calibration function.
-Fortunately, specification of our knowledge in terms of outcome prevalence, c-statistic, and calibration function, that is, defining $\theta=\{\phi,c,h\}$, can, under mild regularity conditions, fully identify $P(\pi,Y)$.
+Fortunately, specification of our knowledge in terms of outcome prevalence, c-statistic, and calibration function (the same components used in Riley’s equations), that is, defining $\theta=\{\phi,c,h\}$, can, under mild regularity conditions, fully identify $P(\pi,Y)$.
 The regularity conditions are as follows:
@@
     $P(p)$ is quantile-identifiable; i.e., any two quantiles of the distribution are sufficient for uniquely identifying it.
-Typical distributions for risks, including Beta, Logit-normal, and Probit-normal satisfy this requirement23.
+Typical distributions for risks, including Beta, Logit-normal, and Probit-normal satisfy this requirement[24].
 The first condition guarantees that the c-statistic relating $\pi$ to $Y$ (which is often reported) is equal to the c-statistic for $P(p)$ (as c-statistic is invariant under monotonical transformation of predictor values).
-The second condition is a requirement for the identifiability of $P(p)$ given its mean (prevalence) and c-statistic23.
-As an example of this identifiability, consider an outcome prevalence of 0.25, c-statistic of 0.75.
-calibration slope of 1.1, and O/E ratio of 0.9.
+The second condition is a requirement for the identifiability of $P(p)$ given its mean (prevalence) and c-statistic[24].
+As an example of this identifiability, consider an outcome prevalence of 0.25, c-statistic of 0.75, calibration slope of 1.1, and O/E ratio of 0.9.
 Assuming predicted risks have a Logit-normal distribution, the calibrated risks will also have Logit-normal distribution.
-The parameters of the latter are uniquely identifiable from $\phi,c$, which resolves to $P(p)\sim\mbox{Logitnorm}(-1.3302,1.0395)$ (the mcmapper R package implements our proposed numerical algorithms for this mapping24).
+The parameters of the latter are uniquely identifiable from $\{\phi,c\}$, which resolves to $P(p)\sim\mbox{Logitnorm}(-1.3302,1.0395)$ (the mcmapper R package implements our proposed numerical algorithms for this mapping[25]).
 As well, there is a 1:1 mapping between the O/E ratio and calibration intercept.
@@
 Given this identifiability, characterizing our prior information involves specifying the joint distribution $P(\phi,c,h)$.
-Ideally, the previous analysis would report both the value and an estimated covariance matrix variance of $\hat{\theta}$, i.e., there would be joint inference about the three elements of theta.
+Ideally, the previous analysis would report both the value and an estimated covariance matrix variance of $\hat{\theta}$; i.e., there would be joint inference about the three elements of $\theta$.
 These would then become the prior mean and covariance matrix of the joint prior for $\theta$.
-In typical practice, however, joint inference on such parameters is not typically reported, but probability distributions for individual components can readily be constructed from existing information.
+In typical practice, however, joint inference on such parameters is not reported, but probability distributions for individual components can readily be constructed from existing information.
 For example, for outcome prevalence, if from a previous study with sample size of $n$, $m$ individuals experience the outcome, our knowledge can be specified as $\mbox{Beta}(m,n-m)$.
-For other components, point estimates and reported CI widths can be used, along an assumed distribution type, to construct distributions.
-Examples include specifying a Log-normal distribution for O/E ratio, Normal distribution for calibration slope, and Beta for c-statistic based on reported point estimates and bounds of 95%CI.
-The accompanying software provides a flexible way of specifying such distributions, accepting both direct specification of distribution parameters, moments, or mean and upper bound of 95%CI.
+For other components, point estimates and reported CI bounds can be used, along an assumed distribution type, to construct distributions.
+Examples include specifying a Log-normal distribution for O/E ratio, Normal distribution for calibration slope, and Beta distribution for c-statistic based on reported point estimates and bounds of 95%CI.
+The accompanying software provides a flexible way of specifying such distributions, accepting specification of distribution parameters, moments, or mean and upper bound of 95%CI.
 Once marginal distributions are constructed, a simple strategy would be to complete the prior specification by imposing a priori independence between the three components.
 An optional strategy, which would be more faithful to the true data generating mechanism, would involve a parametric bootstrap procedure to recover parameters interdependence.
-In this approach, one simulates multiple samples given $theta=\hat{thata}$, and for each sample record the ensuing estimates of $\hat{\theta}$.
+In this approach, one simulates multiple samples given $\theta=\hat{\theta}$, and for each sample records the ensuing estimates of $\hat{\theta}$.
 The empirical correlation matrix of these simulated estimates can then be taken as the prior correlation matrix, instead of simply presuming an uncorrelated prior.
-Appendix 1 provides an algorithmic description of this approach.
 This algorithm in itself quantifies the uncertainty for a population that is exchangeable with the population(s) from which evidence on model performance is collected.
@@
 Independently, if evidence is extracted from a model development study, and there are concerns about the model being overfitted, one can add a negative penalty term to the calibration slope in Step 4b (itself a random variable) representing our knowledge about the degree of overfitting.
-A structured examination of the degree of relatedness between the target population and source population(s) might help the investigator decide on the need for, and extent of, modifications25.
-2.4.3 Sampling from posterior distributions
+A structured examination of the degree of relatedness between the target population and source population(s) might help the investigator decide on the need for, and extent of, modifications[26].
+Sampling from pre-posterior distributions
 All sample size rules require expectation with respect to the distribution of parameters ($\theta$) and future sample ($D_{N}$).
 This naturally invites a Monte Carlo sampling algorithm based on repeated sampling from $P(\theta)$, simulating validation samples from $P(D_{N}|\theta)$, and computing the precision targets and VoI metrics in this sample.
-Draws from the posterior distributions are then processed according to the sample size rules of interest (e.g., average CI widths for the expected CI width criterion, the corresponding quantile for assurance-based metrics, or VoI for NB).
-In addition to this default approach based on simulating $D_{N}$, we propose an approximate two-level closed-form approach for CI width targets that does not involve simulating $D_{N}$.
+Draws from the pre-posterior distributions are then processed according to the sample size rules of interest (e.g., average CI widths for the expected CI width criterion, the corresponding quantile for assurance-based metrics, or VoI for NB).
+In addition to this default approach based on simulating $D_{N}$, we propose an approximate two-step approach for CI width targets that does not involve simulating $D_{N}$.
 This approaches is based on using SE equations twice.
@@
 Table 1 provides an algorithmic description of both approaches.
-Table 1: Bayesian Monte Carlo algorithm for drawing from the posterior distribution of precision targets
+Table 1: Bayesian Monte Carlo algorithm for drawing from the pre-posterior distribution of precision targets
 1.
@@
 (c) Using ${D_{N}}^{(j)}$, construct and record precision targets (CI widths): $ciw^{(j)}=g({D_{N}}^{(j)})$ for each metric of interest.
-*Two-step closed-form method (b) For any metric $\theta_{k}$, specify $P(\hat{\theta}_{k}|\theta^{(j)})$ using method of moments, with the first moment being true value of $\theta_{k}$ from $\theta^{(j)}$, and the second moment being $SE^{2}(\theta_{k})$ (and a choice of distribution type#).
+*Two-step approach (b) For any metric $\theta_{k}$, specify $P(\hat{\theta}_{k}|\theta^{(j)})$ using method of moments, with the first moment being true value of $\theta_{k}$ from $\theta^{(j)}$, and the second moment being $SE^{2}(\theta_{k})$ (and a choice of distribution type#).
 Obtain $\hat{\theta}_{k}$ as a draw from this distribution.
@@
 For VoI analysis, we are not aware of any previous work proposing Optimality Assurance.
-As well, the previously proposed algorithms for validation EVSI were fully Bayesian18.
+As well, the previously proposed algorithms for validation EVSI were fully Bayesian[19].
 Our proposed algorithm for computing Optimality Assurance and EVSI for the mixed Bayesian-likelihood setup is presented in Table 2.
@@
 $N^{(j)}_{fp}=N-N^{(j)}_{+}-N^{(j)}_{tn}$.
-(e) Calculate sample esaimtes of NBs: ${NB}_{D_{N}^{(j)}}(k);k\in\{0,1,2\}$, by plugging in the sample estimates of prevalence, sensitivity, and specificity from ${D_{N}}^{(j)}$ in equation 2.2.
+(e) Calculate sample esaimtes of NBs: ${NB}_{D_{N}^{(j)}}(k);k\in\{0,1,2\}$, by plugging in the sample estimates of prevalence, sensitivity, and specificity from ${D_{N}}^{(j)}$ in equation Common metrics of model performance.
 (f) Identify the winning strategy in the sample: $W^{(j)}=\mbox{argmax}_{k}{NB}_{D_{N}^{(j)}}(k)$.
@@
 The integrals for sensitivity and specificity require numerical methods (with some exceptions, for example, for the Beta distribution).
-2.4.4 Implementation
-The above algorithms are implemented in the bayespmtools package (https://github.com/resplab/bayespmtools), in particular as two functions bpm_valprec() and bpm_valsamp(), for, respectively, computing precision / VoI for a fixed sample size, or determining the sample size corresponding to a set of precision / VoI criteria and sample size rules.
-As input, these functions expect evidence to be parameterized via four probability distributions, one for outcome prevalence, c-statistic, calibration slope, and one of calibration intercept, O/E ratio, or mean calibration.
+Implementation
+The above algorithms are implemented in the accompanying bayespmtools package[27], in particular as two functions bayespm_prec() and bayespm_samp(), for, respectively, computing precision / VoI for a fixed sample size, or determining the sample size corresponding to a set of precision / VoI criteria and sample size rules.
+These functions expect evidence to be parameterized via four probability distributions, one for outcome prevalence, c-statistic, calibration slope, and one of calibration intercept, O/E ratio, or mean calibration.
 Missing correlation among these distributions is optionally (active by default) imputed via the parametric bootstrap method explained earlier.
-The correlation is induced to the sample of draws from marginal draws for these parameters via the methods by Iman and Conover26.
-This method offers the flexibility to the investigator to use different distribution types for each component (as opposed to a single multivariate distribution).
+The correlation is induced to marginal draws for these parameters via the methods by Iman and Conover[28].
+This method offers the flexibility to use different distribution types for each component (as opposed to a single multivariate distribution).
 Both algorithms currently implement Logit-normal, Probit-normal, or Beta for the distribution of calibrated risks.
-bayespm_prec() computes expected CI widths, quantiles of CI widths, as well as optimality assurance and EVSI using the above-mentioned Monte Carlo sampling algorithms.
+For a specified sample size, bayespm_prec() computes expected CI widths, quantiles of CI widths, as well as Optimality Assurance and EVSI using the above-mentioned Monte Carlo sampling algorithms.
 bayespm_samp() solves for minimum $N$ that satisfies any of the requested precision criteria coupled with a sample size rule (e.g., targeting an expected CI width of 0.2 for O/E ratio, or 90% assurance that c-statistic CI width<0.1).
-Because for any $N$, one iteration of the Monte Carlo simulation generates one draw from the posterior distribution of CI widths or NBs, finding the minimum $N$ that satisfies a given sample size rule is a stochastic root-finding problem.
-In our implementation, for targets related to CI widths (expected value and assurance), we use the Robbins-Monro algorithm27, This was motivated by the simplicity of this algorithm which enables simultaneous optimization for all widths targets.
-For VoI and optimality assurance, we note that binomial draws can be vectorized across the entire sample of $\theta$s, thus these calculations can be performed faster than those based on CI Widths.
-We use the Stochastic-Simultaneous Optimistic Optimization algorithm by Valko et al28 (implemented in the OOR R package29).
-3 Case study: Predicting deterioration in hospitalized COVID-19 patients
-The ISARIC 4C model is a risk prediction model for predicting deterioration (need for ventilatory support or critical care, or death) in patients hospitalized due to COVID-19 infection.
-Details of the development and validation on this model are provided in its original publication30.
-In summary, the investigators used electronic hospital records from all 9 health regions of the UK to develop and validate .
-The London region was left out for external validation, and the remaining 8 regions were used in internal-external validation31.
-One at a time, one region was left out, the model was developed in the remaining 7 regions, and its performance was evaluated in the left-out region.
+Because for any $N$, one iteration of the Monte Carlo simulation generates one draw from the pre-posterior distribution of CI widths or NBs, finding the minimum $N$ that satisfies a given sample size rule is a stochastic root-finding problem.
+In our implementation, for targets related to CI widths (expected value and assurance), we use the Robbins-Monro algorithm[29].
+This was motivated by the simplicity of this algorithm which enables simultaneous optimization for all widths targets.
+For Optimality Assurance and VoI metrics, the binomial draws can be vectorized across the entire sample of $\theta$s; thus these calculations can be performed faster than those based on CI widths.
+We use the Stochastic-Simultaneous Optimistic Optimization algorithm by Valko et al[30] (implemented in the OOR R package[31]).
+Case study: Predicting deterioration in hospitalized COVID-19 patients
+The ISARIC 4C model is a risk prediction model for predicting deterioration (need for ventilatory support or critical care, or death) in patients hospitalized due to COVID-19 infection[32].
+The investigators used electronic hospital records from all 9 health regions of the UK to develop and validate this model.
+The London region was left out for external validation, and the remaining 8 regions were used in internal-external validation, where one at a time, one region was left out; the model was fitted using data from all other regions, and its out-of-sample performance was assessed in the left-out region[33].
 Random-effects meta-analysis was used to pool out-of-sample estimates of discrimination and calibration metrics.
 Finally, the investigators used the data from all 8 regions to fit one final model, and evaluated its performance in the London region.
-To use this setup as an informative example, we assume the London region did not participate in this study, and is now interested in conducting a validation study of this model in their region.
+To use this setup as an informative example, we assume the London region did not participate in this study, but is now interested in conducting a validation study of this model in their region.
 That is, we ignore the results reported on the performance of the model in the London region, and consider the internal-external validation results as the information at hand.
-We assume that the performance of the model in the London region can be assumed to be a random draw from the distribution of performance observed across other regions (i.e., it is exchangeable with other regions).
+We assume that the performance of the model in the London region will be a random draw from the distribution of performance observed across other regions (i.e., it is exchangeable with other regions).
 The total development sample size was 70,349 (after removing those with unknown outcome status), with total number of events (deterioration) being 30,316, giving rise to an overall outcome prevalence of 0.428.
-Pooled estimates of the c-statistic, mean calibration, and calibration slope were 0.76 (95%CI 0.75–0.77), -0.01 (95%CI -0.12–0.09), and 0.99 (95%CI 0.97–1.02), respectively.
-3.0.1 Application of Riley approach
-As a baseline comparison, we performed sample size calculations for validating this model using the multi-criteria method proposed by Riley et al3.
-Riley et al used the ISARIC 4C as an example, and for consistency, we use the same targets: confidence interval widths of 0.1 for the c-statistic, 0.22 for O/E ratio, 0.3 for calibration slope.
-The assumed true values of these metrics was taken as the point estimates from the above-mentioned meta-analyses.
+Pooled estimates of the c-statistic, mean calibration (difference between average observed and predicted risks), and calibration slope were 0.76 (95%CI 0.75–0.77), -0.01 (95%CI -0.12–0.09), and 0.99 (95%CI 0.97–1.02), respectively.
+The research question now is: what (minimum) sample size is recommended for a new external validation study in the London region?
+Application of Riley approach
+As a baseline comparison, we performed sample size calculations for validating this model using the multi-criteria method proposed by Riley et al[3].
+Riley et al used the ISARIC 4C as an example, and for consistency, we use the same targets: confidence interval widths of 0.1 for the c-statistic, 0.22 for O/E ratio, and 0.3 for calibration slope.
+The assumed true values of these metrics were taken as the summary (pooled) estimates from the above-mentioned meta-analyses.
 The results indicate a sample size of 1,056 is required, which is dictated by the calibration slope.
 For other components, sample sizes are as follows: O/E ratio: 425, c-statistic: 359.
-3.0.2 Application of Bayesian approach
+Application of Bayesian approach
 Specifying $P(\theta)$: Given our assumption about the exchangeability of populations across UK geographic regions, the predictive distribution of the model performance in a new region represents our uncertainty about the model performance in the London region.
 However, it is crucial to note that pooled estimates and confidence bands reported in the original study are for the average effect and do not represent predictive distributions.
-To obtain predictive distributions, we re-performed the meta-analyses pooling c-statistic, mean calibration (note that while we specify evidence in terms of mean calibration, as is reported by Gupta et al, we target O/E ratio for sample size determination), and calibration slope for the 8 regions in the internal-external validation (we extracted study-specific estimates from digitized figures which enables us to use three significant digits).
+To obtain predictive distributions, we re-performed the meta-analyses pooling c-statistic, mean calibration (note that while we specify evidence in terms of mean calibration, as is reported by Gupta et al, we target O/E ratio for sample size determination), and calibration slope for the 8 regions in the internal-external validation.
 Gupta et al did not perform a meta-analysis on prevalence.
 We therefore performed this additional meta-analysis using region-specific prevalence data from their report.
-Given the large sample sizes, we performed all meta-analyses based on normality assumption on the original scale of each parameter, aside from c-statistic, for which a logit-transformation was performed as recommended by Snell et al32.
-We used method of moments to recover distribution from the mean and SD of the predictive distribution for a new population.
-We specified Beta distributions for prevalence and c-statistic, and Normal distributions for mean calibration and calibration slope.
-This resulted in the distributions for the four parameters presented in Table LABEL:tab:ISARIC_evidence.
+Given the large sample sizes, we performed all meta-analyses based on normality assumption on the original scale of each parameter, aside from c-statistic, for which a logit-transformation was performed as recommended by Snell et al[34].
+We used method of moments to derive distribution parameters from the mean and SD of the predictive distribution for a new population.
+We specified Beta distribution for prevalence, Logit-normal for c-statistic, and Normal for mean calibration and calibration slope.
+This resulted in the distributions for the four parameters presented in Table 3.
 Table 3: Evidence synthesis for the ISARIC study, based on predictive distribution of the internal-external validation results
-Parameter Distribution Mean..SD. CI.bounds
+Parameter Distribution Mean(SD) 95% Credible Interval
 prevalence Beta(119.64,159.91) 0.428 (0.03) 0.371–0.486
@@
 calibration slope Normal(0.9950,0.0237) 0.995 (0.024) 0.949–1.042
-There are some approximations in the above approach that are made in favor of practicality.
-Possible correlations between parameters within each study were not available.We did not model between-study correlations (e.g., via a joint meta-analysis of parameters), after an exploratory analysis (examining Spearman rank correlation coefficients) did not support the presence of strong between-study correlation.
-Target precision / VoI criteria and sample size rules: For comparability, we target the same CI width as in the Riley approach, targeting 95%CI width of, 0.22 for O/E ratio, 0.10 for c-statistic, and 0.30 for calibration slopes.
+We extracted study-specific estimates from digitized figures from the original report which enables us to use three significant digits.
+Matched to the 95% prediction interval from the random-effects meta-analysis - see text.
+We did not model between-study correlations (e.g., via a joint meta-analysis of parameters), after an exploratory analysis (examining Spearman rank correlation coefficients) did not support the presence of strong between-study correlation.
+Target precision / VoI criteria and sample size rules: For comparability, we target the same CI width as in the Riley approach, targeting 95%CI width of 0.22 for O/E ratio, 0.10 for c-statistic, and 0.30 for calibration slopes.
 For these metrics, we apply both the expected value and assurance rules.
-That is, we determine sample sizes that correspond to the expected CI width being equal to our targets (this can be interpreted as a direct Bayesian counterpart of conventional Riley et al’s formula).
-We also demand a 90% assurance to meet or exceed the CI width criteria, representing our preference for obtaining a narrower over wider CI width than the targeted with.
+That is, we determine sample sizes that correspond to the expected CI width being equal to our targets (this can be interpreted as a direct Bayesian counterpart of conventional Riley et al’s targets).
+We also demand a 90% assurance to meet or exceed the CI width criteria, representing our preference for obtaining a narrower over wider CI width.
 As for clinical utility, we demand a 90% assurance on NB at the 0.2 threshold; that is, we desire to be 90% confident that the strategy that will emerge as the best in the validation sample is actually the best strategy in the target population.
 Finally, we investigate the expected gain in NB for each component of the sample size on the EVSI curve.
-Analysis setup: All analyses are based on 10,000 draws from $P(\theta)$, assuming Logit-normal distribution for calibrated risks.
-This assumption is tested in a sensitivity analysis.
-We used the sample-based approach in this analysis.
-Results of the two-level closed-form approach are provided in Appendix 2.
+This is an example of a hybrid use of this framework: CI widths and Optimality Assurance are used as sample size rules, and resulting sample sizes are assessed on in terms of their EVSI values.
+Analysis setup: All analyses were performed in R[35].
+We used the meta package for random-effects meta-analysis of internal-external validation results of ISARIC[36].
+Sample size computations were done using the bayespmtools package, with 10,000 draws from $P(\theta)$, and assuming Logit-normal distribution for calibrated risks.
+We used the sample-based approach for this analysis.
 Results: Table 4 provides the sample size for each component.
-The expected CIW components are close to their frequentist counterparts.
+The expected CI width components are close to their frequentist counterparts.
 The assurance-based targets result in slightly higher sample sizes, as expected.
-The final sample size of 1173 is dictated by the assurance component of the calibration slope.
+The final sample size of 1,181 is dictated by the assurance component of the calibration slope.
 Table 4: Sample size components based onconventional Riley approach (top row) and the Bayesian approach (bottom rows)
@@
 Conventional Riley approach (expected CI widths) 359 425 1056 149
-Bayesian (expected CI width) 360 426 1025
-Bayesian (90% assurance) 397 516 1173 306
+Bayesian (expected CI width) 351 430 1064 NA
+Bayesian (90% assurance) 399 522 1181 306
+Not computed as this framework emphasizes computing assurance and VoI metrics for NB.
 Figure 1 shows how the precision targets change with sample size.
-These are computed from a separate Monte Carlo simulation of size 10,000 (independently of the calculations that determined the sample sizes) and as such act as diagnostic plots of CIW and NB assurance.
-Figure 1: Expected CI widths (a) 90th quantile of CI width (b), optimality assurance (c) curves. Shapes on the line pertain to individual sample size components.
+These are computed from a separate call to bayespm_prec() (independently of the calculations that determined the sample sizes) and as such act as diagnostic plots.
+Figure 1: Expected CI widths (a), 90th quantile of CI width (b), and optimality assurance (c) curves. Shapes on the line pertain to individual sample size components.
 Refer to caption Refer to caption Refer to caption
-(a) Expected value of CI width (b) 90th percentile of CI widths (c) Optimality assurance
-Figure 2 shows the exemplary kernel histograms of the distribution of CI widths (the first three panels) for the smallest ($N=306)$ and largest ($N=1,173$) components of the sample size.
+(a) Expected value of CI widths (b) 90th percentile of CI widths (c) Optimality assurance
+Figure 2 shows the exemplary kernel histograms of the distribution of CI widths (the first three panels) for the smallest ($N=306)$ and largest ($N=1,181$) components of the sample size.
 The last panel demonstrates the distribution of the incremental NB of the model compared with the default strategies (i.e., $NB_{1}-\max(NB_{0}-NB_{2})$).
-With higher sample sizes, the CI widths get both shorter and also get more clustered.
+With higher sample sizes, the CI widths get both shorter and also more clustered.
 The distributions are relatively symmetrical.
 This can explain why the ECIW values are close to their conventional, frequentist counterparts.
-For NB, as the sample estimate of NB is an unbiased estimator, higher sample sizes will result in a narrower distributions.
-The assurance for NB, and for EVSI, what affects the results is the share of the distribution that falls on the left side of zero.
+For NB, as the sample estimate of NB is an unbiased estimator, higher sample sizes will result in a narrower distributions but their location remains the same.
+For Optimality Assurance and EVSI, what affects the results is the share of the distribution that falls on the left side of zero.
 Figure 2: Kernel histograms of CI widths for (a) c-statistic , (b) O/E ratio, and (c) calibration slope.
@@
 Refer to caption Refer to caption
-Turning to VoI analysis, Figure 3 provides the EVSI calculation (with EVSI/EVPI ratio as the secondary Y-axis). Components of sample size calculation are overlaid on the graph.
+Turning to VoI analysis, Figure 3 provides the EVSI calculation (with EVSI/EVPI ratio [rEVSI] as the secondary Y-axis). Components of sample size calculation are overlaid on the graph.
 Figure 3: Expected Value of Sample Information (EVSI) curve. Shapes pertain to individual sample size components
-EVSI quantifies the expected value of a validation sample of a given size to identify the strategy with highest clinical utility
-From this figure, one can (subjectively) realize that the CIW components fall on the ‘diminishing return’ of the EVSI.
-For example, going from the the largest sample size based on criteria other than the calibration slope (N=516) to the sample size dictated by assurance on calibration (N=1,173), a more than doubling of the sample size, is associated with 9.2% expected gain in NB.
-This can justify relaxing the criteria on calibration sloe for the validation study, if the focus is on clinical utility.
+From this figure, one can (subjectively) realize that some of the CIW components fall on the ‘diminishing return’ of the EVSI curve.
+For example, going from the the largest sample size based on criteria other than the calibration slope (N=522) to the sample size dictated by assurance on calibration (N=1,181), a more than doubling of the sample size, is associated with only 8.8% expected gain in NB.
+This can justify relaxing the criteria on calibration slope for the validation study, if the focus is on clinical utility.
 The decision whether to relax the calibration slope criteria can also be examined via the stability of the flexible calibration plots.
-As the desired precision based on the calibration slope is difficult to justify, but often dictates the final sample size required, Riley et al.
-recommend to plot the corresponding variability in calibration curves that might arise for that sample size33.
+As the desired precision based on the calibration slope is difficult to justify, but often dictates the final sample size required, Riley et al recommend plotting the corresponding variability in calibration curves that might arise for that sample size[9].
 In our context, the variability of a calibration curves has two sources: the variability in the true calibration function (due to variability in $\theta$) and the variability due to the finite validation sample (due to variability of $D_{N}$).
@@
 Because in our Bayesian Monte Carlo sampling algorithm, the true calibration function is known within each iteration, we can remove variability due to $\theta$ to quantify the difference between the flexible calibration curve and the true calibration function: for each simulated $D_{N}$, we fit a kernel regression (e.g., LOESS) for $P(Y=1|\pi)$, then remove $h(\pi)$ from the fitted values.
-Figure 4 shows such Bayesian calibration error plots for the final sample sizes including calibration slope criteria (N=1,173) and after removing such criteria (N=516).
+Figure 4 shows such Bayesian calibration error plots for the final sample sizes including calibration slope criteria (N=1,181) and after removing such criteria (N=522).
 Moving to the smaller sample size expectedly increases the spread of calibration errors.
@@
 Again, this can be taken as a further argument in favor of settling on the smaller sample size.
+Putting all these together, the final recommended sample size will be N=522.
 Figure 4: Bayesian calibration error plots - dashed lines are 95% credible bands
 Refer to caption Refer to caption
-4 Discussion
+Discussion
 We proposed a Bayesian framework for sample size considerations when a risk prediction model is to be evaluated (validated) in a target population.
-Compared to the conventional frequentist framework, this approach offers several advantages: it allows investigators to specify their uncertainties about model performance; it enables assurance-type sample size rules that incorporate investigators’ risk preferences; and it facilitates the the use of Value-of-Information analysis when clinical utility of the model is being examined.
+Compared to the conventional frequentist framework, this approach offers several advantages: it allows investigators to specify their uncertainties about model performance; in addition to sample size rules targeting expected CI widths, it enables assurance-type sample size rules that incorporate investigators’ risk preferences; and it facilitates the the use of Value of Information analysis when clinical utility of the model is being examined.
 We proposed characterizing uncertainties around common metrics of model performance, and introduced Monte Carlo sampling algorithms for computing precision and VoI metrics.
-Overall, this framework can provide an objective first step when designing validation studies aimed at quantifying the performance of a pre-specified model in a new population and determining whether it provides clinical utility.
-In this context, this framework helps the investigator explore the trade-off between sample size, precision, and gain in clinical utility.
+We implemented this approach as the bayespmtools R package, with two functions, one for quantifying the anticipated precision / VoI values given a fixed sample size, and one for determining the sample size given pre-specified rules.
+Overall, this framework can provide an objective first step when designing validation studies aimed at quantifying the performance of a pre-specified model in terms of trade-off between sample size, precision, and gain in clinical utility.
 This framework can be used to investigate the anticipated precision or expected gain in clinical utility when the sample size is fixed (based on already collected data).
-Moroever, it can also be used to determine the sample size when original data collection is planned.
+It can also be used to determine the sample size when original data collection is planned.
 A hybrid use, determining the sample size based on certain criteria and examining the resulting precision and VoI for other outcomes, is also possible.
@@
 Among the outputs of this Bayesian approach, the assurance-type rules and VoI metrics are particularly relevant for risk prediction modeling.
-If the estimates of model performance are too uncertain, the adoption of the model might be met with resistance, even if the point estimates are compatible acceptable performance.
-This, combined with the general risk-averse attitude of public funding agencies34, means the costs of not meeting the desired targets is typically more than the benefits of exceeding those targets.
+If the estimates of model performance are too uncertain, the adoption of the model might be met with resistance, even if the point estimates are compatible with acceptable performance.
+This, combined with the general risk-averse attitude of public funding agencies[37], means the costs of not meeting the desired targets is typically more than the benefits of exceeding those targets.
 Assurance-type rules can be used alongside those based on expected values to assuage concerns that the planned study might generate ambiguous results.
-Further, VoI metrics address longstanding criticisms on the irrelevance of inferential statistics when deciding on whether to adopt a health technology (including markers and risk scoring tools) for patient care9.
-These criticism have been recently voiced in the particular case of NB calculations for risk prediction models10.
+Further, VoI metrics address longstanding criticisms on the irrelevance of inferential statistics when deciding on whether to adopt a health technology (including markers and risk scoring tools) for patient care[10].
+These criticism have been recently voiced in the particular case of NB calculations for risk prediction models [11].
 Accordingly, we recommend sample size considerations around NB to move away from confidence intervals and focus on assurance and EVSI.
-In particular, EVSI combines the risk of failing to identify the optimal strategy with the consequences (NB losses) of such a failure, providing a fuller picture of the implications of learning from an empirical study.
-We note that EVSI captures the expected gain in NB per decision.
-When scaled to the population (i.e., multiplied by the expected number of time the decision of interest is to be made), it quantifies the total value of conducting a validation study of a given size in net true or false positive units (an example of such calculations are provided in a previous study18).
+In particular, EVSI combines the risk of failing to identify the optimal strategy with the consequences (NB losses) of such a failure, providing a fuller picture of the implications of learning from an empirical study in terms of clinical utility.
 This provides a ‘value-based’ perspective that can complement the precision-based approach for sample size determination around metrics of discrimination and calibration.
@@
 The application to survival outcomes seems particularly relevant and feasible.
-For such outcomes, model performance needs to be assessed at a time point of interest35.
-The time-dependent equivalents of prevalence and c-statistic are, respectively, the complement of survival probability and time-dependent area under the receiver operating characteristic curve at the time horizon of interest36.
+For such outcomes, model performance needs to be assessed at a time point of interest[38].
+The time-dependent equivalents of prevalence and c-statistic are, respectively, the complement of survival probability and time-dependent area under the Receiver Operating Characteristic curve[39].
 If uncertainty around these metrics are specified, our identifiability conditions can then be employed to map such metrics to the distribution of calibrated risks at this time point.
@@
 We focused on precision and VoI targets that pertain to the population-level performance of the model.
-Sample size consideration can also be approached in terms of uncertainty at the individual level, such as instability at individual-level predictions37.
+Sample size consideration can also be approached in terms of uncertainty at the individual level, such as instability at individual-level predictions[40].
+Further, when contemplating a validation study, often a secondary objective is to update (revise) the model if its performance turns out to be sub-optimal.
+It makes sense to consider targets related to model updating when designing validation studies[41].
 In line with Riley equations, we modeled the calibration function to be linear on the logit scale.
@@
 A more appealing extension would be to model $h()$ non-parametrically.
-This can be done based a non-parametric summaries of calibration curves, such as those based on kernel smoothing (e.g., $ICI$16) or cumulative plots (e.g., $C^{*}$38).
+This can be done based a non-parametric summaries of calibration curves, such as those based on kernel smoothing (e.g., $ICI$[17]) or cumulative plots (e.g., $C^{*}$[42]).
 From an implementation perspective, Bayesian calculations are inherently more complex than their frequentist counterparts.
@@
 These numerical algorithms might struggle for extreme cases.
-In addition, as the Bayesian Monte Carlo sampling produces draws from the posterior distributions of precision targets (e.g., CI widths), sample size determination requires stochastic optimization techniques, for example finding the $N$ that corresponds to the first moment or a quantile of this distribution.
+In addition, as the Bayesian Monte Carlo sampling produces draws from the pre-posterior distributions of precision targets (e.g., CI widths), sample size determination requires stochastic optimization techniques.
 These algorithms demand convergence assessment and if needed repetition with different starting values.
@@
 This practice ignores the fact that predictions for a model trained using a finite sample are inherently uncertain.
-Differences across populations in the relationship between predictors and the outcome, differences in predictor assessment approaches across settings, and variations in outcome ascertainment methods further add to our uncertainties39.
-Fortunately, there have been recent calls on the importance of characterizing and communicating uncertainty around model predictions40.
-These calls encourage investigators to fully document and communicate uncertainty in model coefficients in reports of model development (e.g., reporting the covariance matrix of model coefficients, or reporting the coefficients of models fitted in bootstrapped copies of the original sample)40.
+Differences across populations in the relationship between predictors and the outcome, differences in predictor assessment approaches across settings, and variations in outcome ascertainment methods further add to our uncertainties[43].
+Fortunately, there have been recent calls on the importance of characterizing and communicating uncertainty around model predictions[44].
+These calls encourage investigators to fully document and communicate uncertainty in model coefficients in reports of model development (e.g., reporting the covariance matrix of model coefficients, or reporting the coefficients of models fitted in bootstrapped copies of the original sample)[44].
 These can further facilitate incorporating uncertainties into the design of subsequent studies.
@@
 But, if our existing knowledge is informative enough to warrant investigating the model in a new setting, should it not be used alongside the sample to inform our final judgment?
-Taking our case study as an example, the information on prevalence alone (based on the moments of its distribution) is equal to having learned its value from a sample of 280 individuals - a non-trivial amount of information!
+Taking our case study as an example, the information on prevalence alone (based on the moments of its distribution) is equal to having learned its value from a sample of 281 individuals - a non-trivial amount of information!
 Incorporating such prior knowledge into what we learn from the data will reduce prediction uncertainty, which can affect both patient care (more precise predictions) and the design of empirical studies (requiring smaller samples to achieve the same targets).
@@
 We expect Bayesian approaches towards study design to gain traction as awareness grows regarding the relevance of prediction uncertainty in decision making.
-5 Appendix
-5.1 Appendix 1: Algorithm for parametric bootstrapping
-1.
-For i in 1 to M (size of the Monte Carlo simulation) (a) Generate $D_{n}:=\{(\pi_{i},Y_{i})\}_{i=1}^{n}$, an iid sample of size $n^{\dagger}$ consisting of predicted risks and response values using the point estimates $\hat{\theta}=\{\hat{\phi},\hat{c},\hat{h}\}^{\ddagger}$.
-(b) Estimate $\theta^{(j)}=\{\phi^{(j)},c^{(j)},h^{(j)}\}$ from the above sample.
-2.
-Estimate the correlation coefficients for $\theta$s.
-Generally, correlation coefficients are not sensitive to the size of the sample.
-It is natural to choose $n$ to be the size of the original sample.
-In instances (as in our case study) where elements of $\theta$ are taken from different sources or from the output of meta-analysis (as in our case study), one can determine $n$ based on the effective sample size (e.g., based on uncertainty around prevalence).
-DEtermining the distribution of predicted risks and observed responses from .
-5.2 Appendix 2: Sample size and CI widths for the two-level closed-form method
-Approach c-statistic O/E ratio calibration slope
-Bayesian (expected CI width) 343 420 1044
-Bayesian (90% assurance) 398 516 1173
-Expected CI widths (a) 90th quantile of CI width (b), optimality assurance (c) curves based on the two-level closed-form method. Shapes on the line pertain to individual sample size components.
-Refer to caption Refer to caption
-(a) Expected value of CI width (b) 90th percentile of CI widths